### v1

# Macro Graph of Experts for Billion-Scale Multi-Task Recommendation

Abstract.
Graph-based multi-task learning at billion-scale presents a significant challenge, as different tasks correspond to distinct billion-scale graphs. Traditional multi-task learning methods often neglect these graph structures, relying solely on individual user and item embeddings. However, disregarding graph structures overlooks substantial potential for improving performance. In this paper, we introduce the Macro Graph of Expert (MGOE) framework, the first approach capable of leveraging macro graph embeddings to capture task-specific macro features while modeling the correlations between task-specific experts. Specifically, we propose the concept of a Macro Graph Bottom, which, for the first time, enables multi-task learning models to incorporate graph information effectively. We design the Macro Prediction Tower to dynamically integrate macro knowledge across tasks. MGOE has been deployed at scale, powering multi-task learning for the homepage of a leading billion-scale recommender system. Extensive offline experiments conducted on three public benchmark datasets demonstrate its superiority over state-of-the-art multi-task learning methods, establishing MGOE as a breakthrough in multi-task graph-based recommendation. Furthermore, online A/B tests confirm the superiority of MGOE in billion-scale recommender systems.
multi-task learning, graph neural networks, billion-scale recommender system, click-through rate prediction
copyright: noneccs: Information systems Recommender systemsccs: Information systems Business intelligenceccs: Information systems Online advertising
1. Introduction
Multi-task learning (MTL) lies at the core of nearly all billion-scale recommendation systems (Wang et al., 2023; Caruana, 1997; Kokkinos, 2017), responsible for predicting multiple key tasks such as clicks, likes, and adding to cart (Juan et al., 2016; Zhu et al., 2023). Its goal is to address multiple tasks in real-time to meet diverse user needs. Previous multi-task recommendation models overlook the valuable information embedded within graphs. As shown in Figure 1, their embedding-based multi-task bottoms rely solely on individual user and item embeddings and disregard the rich relational semantics encoded in graph structures, which ultimately results in suboptimal performance. However, incorporating graph structures in MTL presents greater challenges compared to single-task situations due to two primary reasons: 1) online systems have billions of users and items, necessitating efficient and effective handling of this scale, and 2) multiple graph relations are formed by different tasks, requiring appropriate graph management techniques.
The mainstream approach in industrial recommendation typically employs single embeddings for multi-task problems without considering graph structures (Sener and Koltun, 2018; Ruder et al., 2019). ESMM (Ma et al., 2018a) enhances the robustness and improves the overall performance by utilizing the entire exposed sample space for training. MMoE (Ma et al., 2018b) pioneers the use of a basic mixture-of-experts structure, enabling task-specific predictions to learn from other tasks. PLE (Tang et al., 2020) separates task-sharing and task-specific experts to avoid the negative transfer and seesaw phenomenon. STEM (Su et al., 2024) introduces a novel shared and task-specific embeddings paradigm to effectively capture user preferences. Subsequent works focused on designing more sophisticated expert structures, such as multi-layer or task-specialized experts. Other studies explore graph-based multi-task learning; for example, MoGENet (Hu et al., 2022) proposes a multi-channel graph neural networks (GNNs) to model high-order information using user-item bipartite graphs. MMoCEG (Zhang et al., 2023b) introduces graph-based MoE to recognize the commonalities and differences among multiple regions. However, graph-based MTL models often struggle with the high computational burden of GNNs when facing billion-scale graphs.
In general, current multi-task learning models face trade-offs in billion-scale recommender systems:
1. Lack of High-Order Information: Graph neural networks have demonstrated significant performance improvements, but single-embedding MTL approaches cannot benefit from extracting and aggregating high-order information.
2. Computational Complexity: Since each task forms a different graph, managing multiple billion-scale graphs introduces considerable computational overhead for online recommendation systems.
Considering these trade-offs, designing a billion-scale graph neural network structure that addresses multi-task learning scenarios is a promising direction. Recently, MacGNN (Chen et al., 2024) introduced a macro graph structure for single click-through rate (CTR) prediction tasks. However, applying macro structures to multi-task learning presents the following challenges:
1. Multi-task Graph Design: Designing appropriate graph structures to capture and integrate information from multiple tasks is a complex challenge.
2. Macro Expert Design: Developing experts capable of learning graph information to predict different tasks requires careful consideration and architectural innovation.
3. Inter-Macro Merging: Efficiently merging information across macro structures is crucial for effective multi-task learning.
Figure 1. Comparison of Multi-Task Learning Architectures. (a) Embedding-based MTL relies on feature representations. (b) Graph-based MTL utilizes GNNs to capture relational semantics. (c) Macro Graph MTL introduces macro graph bottom to incorporate macro graph information.
To address these challenges, we introduce Macro Graph of Experts (MGOE), the first graph neural network architecture for billion-scale multi-task recommendation that avoids introducing unbearable computational complexity. We present the Macro Task Merging Graph (MTMG), which extracts multi-task information using a single macro graph. We then describe the design of macro graph experts that learn graph information to predict different tasks. Finally, we introduce the Macro Prediction Tower, which aggregates multi-level information to simultaneously predict multiple tasks.
Our contributions can be summarized as follows:
  • We pioneer the use of graph neural networks for billion-scale multi-task recommendation, solving the computational complexity challenges of GNNs in real-world applications.
  • We present the design of MTMG that incorporate multiple different graphs, fundamentally addressing the inability of existing multi-task models to utilize graph information.
  • We provide implementation details for deploying MGOE in a real-world billion-scale recommender system.
  • Extensive offline experiments conducted on three public benchmark datasets verify that MGOE outperforms both embedding-based multi-task learning models and graph-based recommendation models. Besides, online A/B test have confirmed the superiority of MGOE in the billion-scale recommender systems.
2. Preliminaries
2.1. Notations
In this paper, we use $\mathcal{U}$, $\mathcal{I}$, and $\mathcal{T}$ to denote the user set, item set, and task set, respectively. Each task $t\in\mathcal{T}$ is associated with a task-specific interaction matrix ${\mathbf{R}}^{(t)}$. For any user $u\in\mathcal{U}$ and item $i\in\mathcal{I}$, the entry $r^{(t)}_{ui}\in{\mathbf{R}}^{(t)}$ represents the interaction between $u$, $i$, and $t$. Specifically, $r^{(t)}_{ui}=1$ indicates that user $u$ and item $i$ have interacted in task $t$, while $r^{(t)}_{ui}=0$ implies no interaction. For each task $t$, we have a dataset $\mathcal{D}_{t}$, where each element is represented as $(u,i,y_{ui}^{(t)})\in\mathcal{D}_{t}$. Here, the label $y_{ui}^{(t)}\in\{0,1\}$ indicates whether user $u$ interacted (1) or did not interact (0) with item $i$ in task $t$.
2.2. Multi-task Learning
Multi-task learning (MTL) aims to develop the ability to handle multiple related but distinct tasks simultaneously. In e-commerce scenarios, these tasks may involve predicting whether a user will interact with an item by performing a specific behavior, such as clicking on a product, adding it to a cart, or making a purchase. The objective of MTL is to learn a set of functions $\mathcal{F}$, where each function $f_{t}\in\mathcal{F}$ takes user and item features as input to predict the probability of the corresponding task, which is defined as $f_{t}(\cdot):(u,i)\rightarrow\hat{y}_{ui}^{(t)}$, with $\hat{y}_{ui}^{(t)}\in[0,1]$ representing the predicted probability. Typically, MTL employs the binary cross-entropy loss as the loss function for each task $t$, which is expressed as:
$\mathcal{L}_{t}=-\sum_{(u,i,y_{ui}^{(t)})\in\mathcal{D}_{t}}\left[y_{ui}^{(t)}\log(\hat{y}_{ui}^{(t)})+(1-y_{ui}^{(t)})\log(1-\hat{y}_{ui}^{(t)})\right].$
The overall loss function $\mathcal{L}_{\text{MTL}}$ is then defined as:
(1) $\mathcal{L}_{\text{MTL}}=\sum_{t\in\mathcal{T}}\alpha_{t}\mathcal{L}_{t},$
where $\alpha_{t}$ is the parameter to reflect task $t$’s relative importance.
Figure 2. An illustration of the construction of Macro Task Merging Graph and the overview of Macro Graph of Experts.
2.3. MTL Bottom
Embedding-based bottom
The embedding-based MTL bottom serves as the process that transforms feature inputs into fused representations. Typically, for each task, the multi-task bottom takes relevant user micro features ${\bm{e}}_{u}$ and item micro features ${\bm{e}}_{i}$ as input, concatenating them into a vector ${\bm{z}}=[{\bm{e}}_{u};\;{\bm{e}}_{i}]$. This newly formed vector is the input for subsequent expert layers, enabling further feature fusion. Suppose the number of shared experts and task-specific experts is $N_{1}$ and $N_{2}$, respectively. The prediction $\hat{y}_{ui}^{(t)}$ for task $t$ is then expressed as:
(2) $h_{t}\left({\bm{z}}\right)$ $=\sum_{l=1}^{N_{1}}\operatorname{G}_{t,l}^{(1)}\left({\bm{z}}\right)\operatorname{E}_{s,l}\left({\bm{z}}\right)+\sum_{l=1}^{N_{2}}\operatorname{G}_{t,l}^{(2)}\left(\mathbf{z}\right)\operatorname{E}_{t,l}\left({\bm{z}}\right);$
(3) $\hat{y}_{ui}^{(t)}$ $=\sigma\left(\operatorname{MLP}_{t}\left(h_{t}\left({\bm{z}}\right)\right)\right),$
where $\sigma$ denotes the sigmoid function, $\operatorname{E}_{s,l}$ represent the $l$-th shared experts, $\operatorname{E}_{t,l}$ is the $l$-th task-specific experts for task $t$ , $\operatorname{G}_{t,l}^{(1)}(\cdot)$ and $\operatorname{G}_{t,l}^{(2)}(\cdot)$ represent the gating network operations that fuse the outputs from shared experts and task-specific experts, respectively. It is worth noting that embedding-based MTL models separately focus on user and item micro features, neglecting the graph information formed by user-item interactions.
Graph-based bottom
The graph-based bottom typically captures the user-item collaborative connections through graph neural networks (GNNs). Given the input of the $l$-th layer ${\bm{e}}^{l}_{v}$ for a certain user/item node $v$, the output of $(l+1)$-th layer ${\bm{e}}^{l+1}_{v}$ and the prediction $\hat{y}_{ui}^{(t)}$ for task $t$ are represented as:
(4) ${\bm{e}}^{l+1}_{v}$ $=\operatorname{GNN}\left({\bm{e}}^{l}_{v},\sigma\left({\mathbf{W}}_{{\mathcal{G}}_{t}}\cdot\sum_{c\in\mathcal{N}(v)}\eta_{v\leftarrow c}\cdot{\bm{e}}^{l}_{c}\right)\right);$
(5) $\hat{y}_{ui}^{(t)}$ $=\sigma\left(\operatorname{MLP}_{t}\left([{\bm{e}}^{l+1}_{u};\;{\bm{e}}^{l+1}_{i}]\right)\right),$
where ${\mathbf{W}}_{{\mathcal{G}}_{t}}$ is the weight matrix for the task graph ${\mathcal{G}}_{t}$, and $\eta_{v\leftarrow c}$ is implemented using an attention mechanism. Here, $\mathcal{N}(v)$ denotes the set of neighbors of node $v$. However, when dealing with billion-scale task graphs, the graph-based bottom faces significant challenges in computational resources.
3. Methodology
The overall framework of our proposed MGOE is shown in Figure 2. We begin by introducing the macro task merging graph (MTMG) utilized in this paper. Subsequently, we present our proposed MGOE, which is designed to effectively leverage the macro graph information to enhance recommendation performance.
3.1. Macro Task Merging Graph
In this section, we focus on constructing MTMG, which provides crucial macro graph information for the subsequent MGOE. The construction of MTMG primarily involves two core steps: first, obtaining a merging interaction matrix, and second, computing the macro nodes and constructing the macro edges, respectively.
Merging Interaction Matrix
In real-world recommendation scenarios, multiple tasks often coexist (e.g., users may click and purchase an item simultaneously). These task combinations represent the co-existence of various interaction patterns between users and items, reflecting different degrees of user preference.
To accurately capture these preferences, we introduce a scoring function to quantify the importance of such task combinations. For any user $u\in\mathcal{U}$ and item $i\in\mathcal{I}$, the possible task combinations set between them are denoted as:
$\mathbf{C}_{ui}=\{\left(u,i,t_{j_{1}}\right),\left(u,i,t_{j_{2}}\right),\ldots,\left(u,i,t_{j_{k}}\right)\mid t_{j_{1}},t_{j_{2}},\ldots,t_{j_{k}}\in\mathcal{T}\}.$
Let $\mathcal{C}=\{\mathbf{C}_{ui}|u\in\mathcal{U},i\in\mathcal{I}\}$ Once we have defined the importance order of each task, we can obtain a rank function for task combination sets (Zhang et al., 2024a), denoted as $\rho(\cdot):\mathcal{C}\to\mathbb{N}^{+}$. We further define a scoring function $g(\cdot):\mathcal{C}\to\mathbb{R}$, which assigns a merging preference score to each task combination, representing its relative importance. Specifically, the merging preference score $s_{ui}$ between user $u$ and item $i$ is defined as:
(6) $s_{ui}=g\left(\mathbf{C}_{ui}\right)=\begin{cases}\dfrac{[\rho(\mathbf{C}_{ui})+1]^{\beta}-1}{\beta}&,\text{if }\beta\neq 0\\ \ln[\rho(\mathbf{C}_{ui})+1]&,\text{if }\beta=0\end{cases}\,,$
where $\beta$ is a transformation parameter used to adjust the importance of task combination sets. Then, we arrange these values corresponding to user-item pairs to construct the merging interaction matrix ${\mathbf{S}}$:
(7) ${\mathbf{S}}=[s_{ui}]_{m\times n},$
where $s_{ui}$ serves as each entry of the matrix, $m=|\mathcal{U}|$ and $n=|\mathcal{I}|$ denotes the number of users and items respectively. The well-designed scoring function enables systematic quantitative evaluations of different task combinations, ensuring that the constructed MTMG effectively captures the complex interactions between users and items. This, in turn, supports the efficient operation of the MGOE. Following the definition of the scoring function, we proceed to construct the nodes and edges within MTMG.
Macro Node Computation
Computing a macro node involves leveraging the merging preference score computed by Equation 6. Existing approaches treat all nodes in the macro graph uniformly, ignoring the rich semantic relationships in different task combinations. To overcome this limitation, we propose Preference-Based Grouping, a novel method that assigns distinct weights to each micro-node based on its merging preference score. Specifically, we construct macro nodes from two complementary perspectives.
Let ${\mathbf{S}}^{\mathcal{U}}={\mathbf{S}}{\mathbf{S}}^{\top}$ and ${\mathbf{S}}^{\mathcal{I}}={\mathbf{S}}^{\top}{\mathbf{S}}$. Then we can obtain macro nodes of by performing clustering on users and items based on ${\mathbf{S}}^{\mathcal{U}}$ and ${\mathbf{S}}^{\mathcal{I}}$, respectively. The following takes clustering of users as an example. Let $d_{vj}$ denote the $(v,j)$-th element of matrix ${\mathbf{S}}^{\mathcal{U}}$. Subsequently, we assign different weights to different micro nodes. We define a merging preference score baseline $s_{\text{base}}$ for node $v$, such that if the sum of the merging preference score values exceeding $s_{\text{base}}$ surpasses a preset threshold $\delta_{1}$, the node is considered active, and a relatively high weight $w_{\text{high}}$ is assigned. Conversely, if the sum falls below a certain value $\delta_{2}\,\left(\delta_{2}\leq\delta_{1}\right)$, the node is regarded as inactive, and a relatively low weight $w_{\text{low}}$ is assigned. For all other cases, a weight of 1 is assigned. Mathematically, the weight assigned to micro nodes $v$ is defined as:
(8) $w_{v}=\begin{cases}w_{\text{high}}\,,&\sum_{j=1}^{n}\mathbbm{1}_{\{d_{vj}\geq s_{\text{base}}\}}\geq\delta_{1}\\ w_{\text{low}}\,,&\sum_{j=1}^{n}\mathbbm{1}_{\{d_{vj}\geq s_{\text{base}}\}}\leq\delta_{2}\\ 1,&\text{otherwise}\end{cases}\,,$
where $\mathbbm{1}_{\{\cdot\}}$ denotes the indicator function. Then, the centroid update formula of preference-based grouping can be expressed as:
(9) $\bm{\mu}_{k}=\frac{\sum_{v\in\mathcal{V}_{k}}w_{v}\cdot\mathbf{d}_{v}}{\sum_{v\in\mathcal{V}_{k}}w_{v}},$
where $\bm{\mu}_{k}$ is the centroid of cluster $k$, $\mathcal{V}_{k}$ denotes the set of micro-nodes assigned to cluster $k$, $w_{v}$ is the weight associated with node $v$, and $\mathbf{d}_{v}=(d_{v1},d_{v2}\cdots,d_{vn})^{\top}$ is the $v$-th row of ${\mathbf{S}}^{\mathcal{U}}$. The clustering of items can be performed using the same method based on ${\mathbf{S}}^{\mathcal{I}}$. Preference-based grouping ensures that micro-nodes with larger weights exert greater influence on the centroid’s position, thereby making the centroid determination more flexible and better aligned with users’ actual interests and task patterns.
Macro Edge Construction
Within the MTMG, macro edges illustrate the interactions between pairs of macro nodes within defined user or item subgraphs, signifying the task trends within these subgraphs. Unlike micro edges, which link predetermined user and item nodes, macro edges are designed to adaptively represent the varying strengths of connections between two macro nodes, customized for each user or item. Specifically, we formalize the entire macro nodes set in the MTMG as $\widetilde{\mathcal{V}}_{\text{MTMG}}=\left\{\mathcal{V}_{k_{1}},\mathcal{V}_{k_{2}},\ldots,\mathcal{V}_{k_{\widetilde{m}+\widetilde{n}}}\right\}$ with $\widetilde{m}$ and $\widetilde{n}$ denote the number of user clusters and item clusters, respectively. For any node $v$ and its $j^{\text{th}}$-hop neighbors, the weight of the macro edges is denoted by:
(10) $\widetilde{{\mathbf{S}}}_{v;\;\widetilde{p},\,\widetilde{q}}^{(j)}=\sum_{a\in{\mathcal{V}}_{v;\;\widetilde{p}}^{(j-1)}}\sum_{b\in{\mathcal{V}}_{v;\;\widetilde{q}}^{(j)}}s_{ab}\,,$
where $\mathcal{V}_{v;\;\widetilde{p}}^{(j-1)}$ = $\mathcal{V}_{v;\;\widetilde{p}}\ \cap\ \mathcal{N}_{v}^{(j-1)}$ denotes the macro nodes related to node $v$ within its $(j-1)^{\text{th}}$-hop neighbors and $\mathcal{V}_{v;\;\widetilde{q}}^{(j)}$ = $\mathcal{V}_{v;\;\widetilde{q}}\ \cap\ \mathcal{N}_{v}^{(j)}$ represent the macro nodes related to node $v$ within its $j^{\text{th}}$-hop neighbors. By merging different task combinations, MTMG provides a macro-level view of user-item interactions. This approach allows us to observe overarching trends and structures that might not be evident from individual interactions alone.
Macro Task Merging Graph Definition
Our proposed MTMG can be formally defined as $\widetilde{{\mathcal{G}}}_{\text{MTMG}}=\left(\widetilde{\mathcal{U}},\widetilde{\mathcal{I}},\widetilde{{\mathbf{S}}}\right)$, where $\widetilde{\mathcal{U}}$ represents the set of macro user nodes and $\widetilde{\mathcal{I}}$ denotes set of the macro item nodes. Each macro node is associated with a trainable macro embedding $\widetilde{\textbf{E}}$, and its macro neighbors are represented as $\widetilde{\mathcal{N}}$. MTMG captures task patterns from a macro perspective and provides macro graph information for the subsequent macro graph bottom.
3.2. Macro Graph Bottom
In this paper, we propose the macro graph bottom (MGB) with a mixture of macro graph experts (MGE) to exploit rich information from a macro perspective. In MGB, for each task $t$ and a certain node $v$, we assign task-specific macro embeddings to adequately capture the unique macro characteristics:
(11) $\widetilde{{\bm{e}}_{v}^{t}}=\operatorname{Lookup}\left(v,\widetilde{{\mathbf{E}}}_{t}\right)$
where $\widetilde{{\bm{e}}_{v}^{t}}$ represent the macro embeddings. In addition to its own micro features, MGB applies the tailored MGEs to extract information from macro neighbors, achieving a more comprehensive understanding of the relationships and interactions between entities in MTMG from a macro perspective. For a given user-item pair from $\mathcal{D}_{t}$ in MTMG, the input of MGB is concatenated by its micro embedding and macro neighbors embedding:
(12) ${\bm{z}}_{t}=[{\bm{e}}_{u};\;{\bm{e}}_{i};\;\widetilde{{\mathbf{E}}}_{t}(\widetilde{\mathcal{N}}_{u});\;\widetilde{{\mathbf{E}}}_{t}(\widetilde{\mathcal{N}}_{i})],$
where ${\bm{e}}_{u}$ and ${\bm{e}}_{i}$ represents the original micro features and $\widetilde{{\mathbf{E}}}_{b}(\widetilde{\mathcal{N}}_{v})$ denotes the corresponding macro embeddings of $v$’s macro neighbors, and the task $t$ output of MGB is illustrated as follow:
(13) $\operatorname{MGB}\left(\mathcal{D}_{t},\widetilde{{\mathbf{E}}},\widetilde{\mathcal{N}}\right)=\sum_{{t^{\prime}}\in\mathcal{T}}\operatorname{Gate}_{t^{\prime}}\left({\bm{z}}_{t}\right)\cdot\operatorname{MGE}_{t^{\prime}}\left({\bm{z}}_{t^{\prime}}\right),$
where $\operatorname{Gate}_{t^{\prime}}(\cdot)$ denotes the gating network operation, and $\operatorname{MGE}_{t^{\prime}}(\cdot)$ is defined as Equation 17. $\operatorname{MGB}(\cdot)$ integrates user and item embeddings with their corresponding macro neighbor embeddings, resulting in a more informative representation that reflects the overall macro structure.
3.3. Macro Graph Expert
Embedding-based multi-task experts, implemented using an MLP, face limitations in effectively leveraging the macro information in MTMG. We introduce a tailored MGE designed to exploit macro information efficiently. Specifically, we compute macro weights to determine the preference of the target node $v$ with respect to its macro neighbors. These macro weights are derived from the connected macro edges. Furthermore, different task patterns can cause significant disparities in macro edge weights, which we address using logarithmic smoothing. Formally, for a target node $v$ with its $j^{\text{th}}$-hop macro neighbor $\widetilde{p}$, we apply logarithmic smoothing with a parameter $\gamma$ to measure the importance of these macro neighbors. The macro weight $w_{v;\;\widetilde{p}}^{(j)}$ is computed as follows:
(14) $w_{v;\;\widetilde{p}}^{(j)}=\operatorname{softmax}\left(\log\left(\sum_{\widetilde{q}\in\widetilde{\mathcal{N}}_{v}^{(j-1)}}\widetilde{{\mathbf{S}}}_{v;\;\widetilde{p},\,\widetilde{q}}^{(j)}+1\right),\gamma\right),$
where $\gamma$ is a hyperparameter controlling the degree of smoothing (Chen et al., 2023). To aggregate macro embeddings efficiently, we employ an attention mechanism (Vaswani et al., 2017) instead of computationally expensive graph convolutional neural networks. The aggregated macro embedding $\widetilde{{\mathbf{Z}}}_{v,\widetilde{p}}$ is given by:
(15) ${\mathbf{M}}_{\text{sim}}$ $=\frac{({\mathbf{Q}}_{v}\cdot\widetilde{{\mathbf{E}}}_{\widetilde{p}})\cdot\left({\mathbf{K}}_{v}\cdot{\mathbf{E}}_{v}\right)^{\top}}{\sum_{\widetilde{n}\in\widetilde{\mathcal{N}}_{v}^{(j)}\setminus\{\widetilde{p}\}}({\mathbf{Q}}_{v}\cdot\widetilde{{\mathbf{E}}}_{\widetilde{n}})\cdot\left({\mathbf{K}}_{v}\cdot{\mathbf{E}}_{v}\right)^{\top}};$
(16) $\widetilde{{\mathbf{Z}}}_{v,\widetilde{p}}$ $=\operatorname{softmax}(\frac{{\mathbf{M}}_{\text{sim}}}{\sqrt{d}})\cdot\left({\mathbf{V}}_{v}\cdot\widetilde{{\mathbf{E}}}_{\widetilde{p}}\right),$
where ${\mathbf{Q}}_{v},{\mathbf{K}}_{v}$ and ${\mathbf{V}}_{v}$ are learnable query, key, and value matrices for node $v$, and ${\mathbf{M}}_{\text{sim}}$ reflects the similarity between target node $v$ and macro nodes $\widetilde{p}$. Finally, the fusion of macro neighbor representations of task $t$ is expressed as:
(17) $\operatorname{MGE}_{t}({\bm{z}}_{t})=\sum_{\widetilde{p}\in\widetilde{\mathcal{N}}_{v}^{(j)}}w_{v,\widetilde{p}}\cdot\widetilde{{\mathbf{Z}}}_{v,\,\widetilde{p}}\,,$
where $\operatorname{MGE}_{t}({\bm{z}}_{t})$ represents the final output of macro graph expert.
3.4. Macro Prediction Tower
Macro and Micro Fusion
To capture the detailed local interactions provided by micro neighbors and the broader context offered by macro neighbors, we introduce the macro and micro fusion to aggregate information from both perspectives. Specifically, for a given target node $v$ with its numerous neighbors $\mathcal{N}_{v}$ related to task $t$, we sum its micro neighbors. Then, the micro neighbors representation ${\bm{e}}_{v;\;\text{nbr}}^{t}$ can be expressed as:
(18) ${\bm{e}}_{v;\;\text{nbr}}^{t}=\sum_{c\in\mathcal{N}_{v}}\langle{\bm{e}}^{t}_{c},{\bm{e}}_{v}^{t}\rangle\cdot{\bm{e}}^{t}_{c},$
where $\langle{\bm{e}}^{t}_{c},{\bm{e}}_{v}^{t}\rangle$ indicates the weight between the target node and its micro neighbors. Then, we mix the macro and micro informative representations through a multi-gate expert layer:
(19) ${\bm{e}}_{\text{cat}}^{t}$ $=[{\bm{e}}_{u};\;{\bm{e}}_{i};\;\operatorname{MGB}({\bm{z}}_{t});\;{\bm{e}}_{u;\;\text{nbr}}^{t};\;{\bm{e}}_{i;\;\text{nbr}}^{t}];$
(20) $x^{t}$ $=h_{t}\left({\bm{e}}_{\text{cat}}^{t}\right),$
where ${\bm{e}}_{\text{cat}}$ represents the concatenated embedding, $h_{t}(\cdot)$ is defined as Equation 2, and $x^{t}$ denotes the informative fused output from both the macro and micro levels.
To exploit possible connections between tasks and fully utilize a user’s macro preferences for sparse tasks, we replace the traditional multi-task tower with a macro correlated layer. Specifically, we use a macro cascading readout layer to explore the relationships across different tasks and apply macro task adjustment to align the recommendations with the user’s macro preferences.
Macro Cascading Readout Module
Traditional multi-task recommendation models (Ma et al., 2018b; Tang et al., 2020) obtain results in a parallel manner, which ignores the possible connections between different tasks. To address this, we introduce the macro cascading readout module that generates the final output in a cascaded fashion. Specifically, the output of the previous task is fed into the next prediction tower as additional input. Suppose $\hat{y}_{ui}^{(t_{j})}$ is the output of the previous task, then $\hat{y}_{ui}^{(t_{j})}$ can be expressed as:
(21) $\hat{y}_{ui}^{(t_{j})}=\begin{cases}\sigma(\operatorname{MLP}(x^{t_{1}}))\,,&t_{j}=t_{1}\\ \sigma(\operatorname{MLP}(x^{t_{j}})+w_{t_{j-1}}\cdot\hat{y}_{t_{j-1}})\,,&t_{j}\in\mathcal{T}\setminus\{t_{1}\}\\ \end{cases}\,,$
where $w_{t_{j-1}}$ is a trainable parameter that controls the strength of learning the correlation between different tasks.
Macro Task Adjustment
With the goal of recommending more precisely to the user’s individual preferences and enhancing the model’s ability to predict sparse tasks, we propose macro task adjustment to guide the final prediction. First, based on MTMG, we construct a user macro task lookup table according to different task interactions. Specifically, we associate various user’s task interactions with the corresponding macro nodes. For a certain user $u$, the final prediction $\hat{\mathrm{Y}}$ can be expressed as:
(22) $\hat{\mathrm{Y}}={\mathbf{W}}_{u}\cdot{\mathbf{Y}}_{\text{cat}}=\sigma(T_{u})\cdot[\hat{y}_{ui}^{(t_{1})};\;\hat{y}_{ui}^{(t_{2})};\;\ldots]\,,$
where $T_{u}$ is the users’ macro task lookup table, ${\mathbf{W}}_{u}$ represents the macro task adjustment weight corresponding to the user in relation to the target item’s macro node.
3.5. Model Training
We utilize binary cross-entropy as the loss function during the model’s training. Specifically, with the loss function $\mathcal{L}_{t}$ for a certain task $t$, the objective of MTL is given by:
(23) $\mathcal{L}_{\text{MTL}}=\sum_{t\in\mathcal{T}}\alpha_{t}\mathcal{L}_{t}.$
Then, the overall objective function for MGOE is formulated as:
(24) $\mathcal{L}_{\text{MGOE}}=\mathcal{L}_{\text{MTL}}+\lambda_{1}\cdot\|\mathbf{\theta}\|_{1}+\lambda_{2}\cdot\|\mathbf{\theta}\|_{2}^{2}\,,$
where $\lambda_{1}\cdot\|\mathbf{\theta}\|_{1}$ and $\lambda_{2}\cdot\|\mathbf{\theta}\|_{2}^{2}$ represent the $\ell_{1}$ and $\ell_{2}$ regularization terms, respectively, included to prevent overfitting.
4. Experiments
Table 1. Multi-task recommendation performance comparison results. The best and second-best results in each row are highlighted in bold font and underlined, respectively.
Models MMoE PLE ESMM AITM STEM MoME MoGENet MMoCEG POGCN MacGNN MGOE Improv.(%)
Taobao Click AUC 0.5580 0.5578 0.5577 0.5584 0.5561 0.5691 0.5784 0.5720 0.5841 0.5891 0.6158 4.53%
Logloss 0.3730 0.3723 0.3749 0.3922 0.3765 0.4415 0.4897 0.4134 0.3667 0.3641 0.3609 0.88%
Favor AUC 0.5672 0.5663 0.6034 0.5688 0.6192 0.6481 0.6412 0.6250 0.5693 0.5967 0.8016 23.68%
Logloss 0.1525 0.1519 0.1496 0.1628 0.1580 0.1737 0.2211 0.1607 0.1547 0.1497 0.1301 13.03%
Cart AUC 0.5459 0.5469 0.5660 0.5493 0.5666 0.5744 0.5823 0.5679 0.5714 0.5837 0.6652 13.96%
Logloss 0.2470 0.2472 0.2471 0.2666 0.2534 0.2889 0.3333 0.2679 0.2458 0.2424 0.2351 3.01%
Buy AUC 0.5560 0.5550 0.5549 0.5686 0.5676 0.5836 0.6286 0.5971 0.6142 0.6927 0.7098 2.47%
Logloss 0.0966 0.0963 0.0961 0.1012 0.1038 0.1158 0.1356 0.1012 0.0972 0.0930 0.0907 2.47%
Average AUC 0.5567 0.5565 0.5705 0.5612 0.5773 0.5937 0.6076 0.5905 0.5847 0.6155 0.6981 13.42%
Logloss 0.2172 0.2169 0.2169 0.2307 0.2229 0.2549 0.2949 0.2358 0.2161 0.2120 0.2042 3.68%
QB-video Click AUC 0.7891 0.7914 0.7934 0.7936 0.7910 0.8289 0.7786 0.8386 0.8402 0.8415 0.8474 0.70%
Logloss 0.4569 0.4560 0.4547 0.4524 0.4593 0.4337 0.4731 0.4199 0.4192 0.4191 0.4047 3.44%
Like AUC 0.6468 0.6328 0.7361 0.7547 0.7539 0.7250 0.5787 0.7665 0.6673 0.6786 0.8533 11.32%
Logloss 0.0403 0.0400 0.0350 0.0358 0.0408 0.0368 0.0727 0.0357 0.0428 0.0396 0.0300 14.29%
Share AUC 0.5125 0.5016 0.5726 0.5122 0.6403 0.5385 0.5524 0.6110 0.5270 0.6188 0.6925 8.15%
Logloss 0.0088 0.0086 0.0081 0.0167 0.0107 0.0081 0.0243 0.0133 0.0143 0.0080 0.0078 2.50%
Follow AUC 0.6128 0.6067 0.6606 0.5000 0.6792 0.6299 0.6277 0.6915 0.6336 0.6504 0.7424 7.36%
Logloss 0.0073 0.0076 0.0070 0.0131 0.0076 0.0072 0.0230 0.0080 0.0131 0.0071 0.0069 1.43%
Average AUC 0.6403 0.6331 0.6906 0.6401 0.7161 0.6805 0.6343 0.7269 0.6670 0.6973 0.7839 7.84%
Logloss 0.1283 0.1281 0.1262 0.1295 0.1296 0.1215 0.1483 0.1192 0.1224 0.1185 0.1123 5.23%
QK-article Read AUC 0.7089 0.7070 0.7124 0.7080 0.7249 0.7415 0.7732 0.7667 0.7387 0.7911 0.8006 1.20%
Logloss 0.1421 0.1514 0.1164 0.1284 0.1139 0.1077 0.1128 0.1174 0.1119 0.0894 0.0879 1.68%
Like AUC 0.7969 0.7937 0.8118 0.7964 0.8165 0.8028 0.8036 0.8338 0.7959 0.8701 0.9327 7.19%
Logloss 0.0865 0.0936 0.0721 0.0883 0.0719 0.0775 0.0815 0.0830 0.0795 0.0658 0.0519 21.12%
Favor AUC 0.8222 0.8228 0.8344 0.8226 0.8286 0.8315 0.8171 0.8516 0.7911 0.8891 0.9294 4.53%
Logloss 0.0393 0.0403 0.0323 0.0409 0.0329 0.0360 0.0389 0.0349 0.0392 0.0309 0.0256 17.15%
Follow AUC 0.7329 0.7326 0.6664 0.7427 0.7345 0.7404 0.6381 0.7620 0.7054 0.8075 0.8174 1.23%
Logloss 0.0088 0.0087 0.0077 0.0089 0.0073 0.0077 0.0157 0.0076 0.0126 0.0072 0.0070 2.78%
Average AUC 0.7652 0.7640 0.7562 0.7674 0.7761 0.7790 0.7580 0.8035 0.7578 0.8394 0.8700 3.65%
Logloss 0.0692 0.0735 0.0571 0.0666 0.0565 0.0572 0.0622 0.0607 0.0608 0.0483 0.0431 10.77%
In this section, we conduct comprehensive experiments on three real-world datasets, aiming to answer the following research questions. RQ1: How does MGOE perform compared to current state-of-the-art multi-task models? RQ2: What are the effects of the different components of MGOE on the overall performance? RQ3: In what way do the hyper-parameters in MGOE affect the overall performance? RQ4: How efficient is the proposed MGOE? RQ5: How does MGOE perform on billion-scale recommendation platforms?
4.1. Experimental Setup
4.1.1. Datasets
We conduct extensive experiments on three real-world benchmark datasets: Taobao (Zhu et al., 2018), QB-video (Yuan et al., 2022) and QK-article (Yuan et al., 2022). These datasets include a variety of scenarios, including e-commerce and content-based recommendations, to verify the generality and effectiveness of MGOE in the real-world recommendation. The detailed description and the statistics of these datasets is illustrated in B.1.
4.1.2. Compared Baselines
To verify the effectiveness of MGOE, we compare with ten relevant representative multi-task models, which can be roughly divided into two categories. (i)Traditional Multi-task Models: MMOE, PLE, ESMM, AITM, STEM, MoME. (ii) Graph-based Methods: MoGENet, MMoCEG, POGCN, MacGNN. We leave the details of these baseline models in B.2.
4.1.3. Implementation Details
All baselines and MGOE is implemented based on Pytorch. In the offline experiments, the embedding size fix to 10 and use Xavier (Glorot and Bengio, 2010) for initialization for all models. For MGOE, we set the batch size to 1024 and use Adam (Kingma, 2014) for optimization. The learning rate is searched form {1e-2, 5e-2, 1e-3, 5e-3} and the regularization term is searched form {1e-4, 5e-5, 1e-5}.
4.1.4. Evaluation Metrics
We adopt two widely used multi-task evaluation metrics, AUC (Fawcett, 2006) and Logloss (Chen et al., 2016), for a comprehensive evaluation of the models. Higher values of AUC indicate better prediction performance of the model, while the lower values of Logloss imply better prediction accuracy. Higher values of AUC indicate better CTR prediction performance of the model, while the lower values of Logloss indicate better CTR prediction performance. We run the experiments five times with different random seeds and take the average result to avoid extreme cases.
4.2. Main Results (RQ1)
In this subsection, we compare our tailored MGOE with ten state-of-the-art baseline models on three experimental datasets. The comparison results on the AUC and Logloss metrics are reported in Table 1. From the result, we can have the following observations:
MGOE can achieve significant improvements over state-of-the-art methods across all types of tasks on all experimental datasets. From Table 1, we observe that the customized MGOE attains the highest average AUC values and the lowest average Logloss values across all task types, outperforming both traditional multi-task models and graph-based methods. Specifically, in terms of the AUC metric, MGOE surpasses the best baseline with average AUC improvements of 13.42% on Taobao, 7.84% on QB-video, and 3.65% on QK-article. Regarding the logloss metric, MGOE achieves average effective gains of 3.68%, 5.23%, and 10.77% respectively. These comparison results validate that MGOE can produce more reasonable multi-task recommendation outcomes.
Introducing graph information can effectively improve the performance of the models. Based on the results, we can observe that graph-based models, like MMoCEG and MacGNN, generally yield better results compared to traditional multi-task models. This implies that traditional multi-task models treat users and items in isolation, overlooking the relationship between them. Conversely, integrating the information of the user-item bipartite graph can mirror the more profound interactions between users and items. This is of great significance for enhancing the overall recommendation performance.
Introducing graph information from the micro level can not fully utilize the task patterns. By comparing the micro-graph based methods and macro-graph models in Table 1, it can be observed that macro-graph models are capable of achieving relatively superior performance in multi-task recommendation compared to micro-graph ones. This indicates that the traditional microscopic graph fails to effectively translate the diverse tasks of users towards items into enhanced model performance. Consequently, there is a need to integrate these user task patterns in a novel manner and consider these user and item nodes from a macroscopic perspective, which serves as the impetus for proposing MTMG.
Table 2. Ablation study results between MGOE with its three variants on Taobao and QB-video.
Variant Avg. AUC ($\uparrow$) Avg. Logloss ($\downarrow$)
Taobao MGOE 0.6981 0.2042
w/o preference 0.6912 0.2054
w/o adjustment 0.6854 0.2059
w/o cascading 0.6907 0.2054
QB-video MGOE 0.7839 0.1123
w/o preference 0.7806 0.1130
w/o adjustment 0.7659 0.1211
w/o cascading 0.7716 0.1164
4.3. Ablation Study (RQ2)
In order to verify the effectiveness of key designed components in MGOE, we conduct the ablation study and compare MGOE with its three variants: (1) w/o preference removes the preference-based grouping module when constructing MTMG, which ignores the task combination. (2) w/o adjustment excludes the macro task adjustment to overlook users’ macro interest. (3) w/o cascading, which the final output of MGOE is in a parallel fashion.
We conduct ablation studies on Taobao and QB-video datasets. The average of AUC and Logloss computed over all tasks is employed as the evaluation indicator. The results are presented in Table 2. By analyzing this table, we arrive at the following observations:
  • Distinguishing the micro nodes helps form a better macro graph. Ignoring the task relations between users and items decreases the performance of MGOE. This indicates MTMG is a more suitable macro graph for multi-task recommendation.
  • The macro prediction tower facilitates the macro information into more accurate recommendations. Removing the macro cascading readout module leads to a decline in performance. This indicates the necessity of representing the transfer of macro knowledge across different tasks. When the macro task adjustment is removed, the performance deteriorates, as the macro nodes preserve a wealth of users’ preference information.
Figure 3. Parameter study of $\gamma$ and $\widetilde{n}$ on Taobao and QB-video.
Figure 4. Parameter study of $w_{v}$ on Taobao and QB-video.
4.4. Hyper-parameter Study (RQ3)
We investigate three key hyperparameters in MGOE: (1) the temperature parameter $\gamma$ for weighting macro nodes; (2) the number of macro nodes in MTMG, denoted as $\widetilde{n}$; (3) the weight $w_{v}$ in preference-based grouping. To explore the impact of these hyperparameters on MGOE, we conduct an analysis by varying the settings of the key hyperparameters on the Taobao and QB-video datasets. While examining $\gamma$, $\widetilde{n}$, and $w_{v}$, we keep all other hyperparameters fixed. The results are shown in Figure 3 and Figure 4.
Effect of Temperature Parameter. We evaluate the impact of the temperature parameter $\gamma$ over the range from 0.1 to 1.9 with an increment of 0.2. The results are shown in the top half of Figure 3. From the line chart, we can observe that, in the Taobao dataset, as $\gamma$ increases, the model performance initially improves and then declines. The performance reaches its peak when $\gamma=1.1$. In the QB-video dataset, MGOE performs best when $\gamma=1.5$, and then declines with the increase of $\gamma$.
Effect of Macro Node Number. We vary the number of macro nodes from 5 to 30 in steps of 5. From the results presented in the bottom half of Figure 3, we can observe that an overly small number of macro nodes tends to lead to overly coarse segmentation, resulting in poor outcomes. Additionally, selecting a relatively appropriate number of macro nodes can yield satisfactory performance. For both Taobao and QB-video, when the number of macro nodes reaches 20, MGOE achieves the best performance.
Effect of Preference-Based Grouping Weight. We set the value of $w_{\text{high}}$ to either 2 or 1, and $w_{\text{low}}$ to either 1 or 0.5. Thus, we can represent the four combinations as follows: $w_{1}$=(2, 0.5), $w_{2}$=(2, 1), $w_{3}$=(1, 0.5) and $w_{4}$=(1, 1), where the first value corresponds to $w_{\text{high}}$ and the second to $w_{\text{low}}$. From the results presented in Figure 4, we can observe that in both datasets, assigning larger weights to active micro nodes and relatively smaller weight to non-active micro nodes generally leads to better performance.
Figure 5. Efficiency study of the model inference time.
4.5. Efficiency Study (RQ4)
Since multi-task recommendation models are often deployed in real-time scenarios where multiple tasks need to be handled simultaneously, the computational efficiency of these models is a crucial evaluation factor (Wang et al., 2019b). Hence, to verify the efficiency of our proposed MGOE, we compare the average inference time between MGOE and several representative and graph-based multi-task baselines: expert-based model MMoE and task-specific embedding scheme STEM, graph-based method POGCN, multi-channels paradigm MoGENet and graph MOE-based MMoCEG.
The comparison results are presented in Figure 5. From the figure, we have the following observations: (i) Our proposed MGOE is nearly as efficient as the simplest MMoE model and on par with STEM in terms of efficiency. While maintaining the same inference time as these models, our model incorporates high-order graph information, which significantly enhances the performance of multi-task recommendation. (ii) Our model has a much shorter inference time compared with traditional graph-based multi-task models. MoGENet requires almost four times the inference time of our model, indicating that our approach to integrating high-order graph information is far more efficient, enabling faster and more responsive multi-task recommendations, which is of great significance for real-time applications in large-scale recommendation systems.
4.6. Online Evaluation (RQ5)
We conducted an online A/B test on Alibaba platform. In this experiment, our model served as a rank model, replacing the existing online best-performed ranking models—MMOE and MacGNN. MacGNN mainly considers the clicking task, MMOE and MGOE consider “clicking”, “adding to cart”, “favoring”, and “buying”. Table 3 presents the average relative performance variation over two weeks for about 0.5 billion users and 1.2 billion items.
Table 3. Results of online A/B tests in the Alibaba platform.
A/B Test PCTR UCTR CVR GMV StayTime
v.s. MMOE +2.16% +1.63% + 5.88% + 16.46% +4.12%
v.s. MacGNN +0.93% +0.75% +3.69% +7.74% +3.26%
Compared to MMOE, firstly, MGOE demonstrates a performance improvement of 2.16% for PCTR, 1.63% for UCTR, 5.88% for CVR, and 16.46% for GMV, suggesting that our model boosts users’ inclination to interact with items and turn them into purchases. Secondly, the StayTime increases by 4.12%. This indicates that, despite both models considering a wide range of user actions, by incorporating their comprehensive macro interests, our model is more effective in attracting user clicks and driving business revenue.
Compared to MacGNN, which has a more narrow focus on clicking task, our model still shows favorable performance. The PCTR increased by 0.93%, the UCTR rose by 0.75%, the CVR increased by 3.69%, the GMV grew by 7.74%, and the user StayTime increased by 3.26%. This suggests that the broader consideration of user actions in MGOE can lead to better overall business outcomes, even when compared to a model specialized in a single task.
Overall, in the online A/B test on the Alibaba platform, all the results show that MTMG and the equipped MGOE are more suitable than the state-of-the-art online multi-task recommendation models.
5. Conclusion
The introduction of the Macro Task Merging Graph (MTMG) and Macro Graph of Experts (MGOE) not only significantly advances the field of multi-task graph-based recommendation but also has practical implications for improving the performance of large-scale recommender systems. MTMG merges multiple billion-scale graphs into a unified macro structure, effectively consolidating complex graph data. MGOE leverages macro graph embeddings with macro graph experts, which provides a novel approach for integrating macro graph information and enables more efficient information extraction. Extensive offline experiments and online A/B tests demonstrate the superiority of our model.
Appendix A RELATED WORK
A.1. Multi-task Learning
Multi-task learning aims to enhance the model’s generalization ability by training multiple tasks concurrently within the same model. The core of this method lies in the knowledge that can be shared among different tasks, enabling them to benefit from one another. In the early days of deep learning, the hard sharing mechanism of parameters was commonly employed. It could be applied to all hidden layers for all tasks, while maintaining task-specific output layers. The parameter soft sharing mechanism, on the other hand, allows each task to have its own model and parameters, thereby achieving better performance. The cross-stitch network utilizes the cross-stitch unit to take a task-specific network and linearly combine it with the previous output by leveraging knowledge learned from other tasks. However, in recommendation systems, parameter sharing between tasks that have poor correlation can lead to a seesaw phenomenon. To solve the issues of data sparsity (DS) and sample selection bias (SSB), ESSM (Ma et al., 2018a) simultaneously optimizes two related tasks in a sequential manner. MOE (Jacobs et al., 1991) employs a gated network to transform the original parameters shared by all samples into multiple sets of parameters, where each set is called an expert. MMOE (Ma et al., 2018b) proposes a multi-expert network to address the multi-task learning problem, with the hope that the model could automatically learn to assign expert parameters based on the relationship of the underlying task. PLE (Tang et al., 2020) divides the model parameters into a private part and a shared part, which improves the robustness of multi-task learning and mitigates the negative impact between private knowledge and shared knowledge.
A.2. Graph Learning for Recommendation
In recent years, an increasing number of studies have attempted to integrate graphs into recommendation systems in order to enhance performance. NGCF (Wang et al., 2019a) typically incorporates higher-order graph information into traditional collaborative filtering. To achieve superior performance, LightGCN (He et al., 2020) removes non-linear operations and feature transformations, which are found to be redundant through experimentation. These methods have been widely adopted in industrial item recall scenarios. However, due to their long inference times, these methods are difficult to apply directly to multi-task learning. Subsequently, some approaches attempt to incorporate graph information into the CTR prediction scenario. MoGENet (Hu et al., 2022) proposes a novel multi-channel graph neural network to jointly model high-order information using the user-item bipartite graph. MMoCEG (Zhang et al., 2023b) utilizes a graph-based structure to share information across regions and employs contrastive learning to prevent expert representation degeneration. GLSM (Sun et al., 2022) employs an adaptive fusion mechanism to fuse information from long-term and short-term behaviors. GMT (Min et al., 2022) effectively integrates different types of interactions between local neighborhood nodes from both explicit and implicit perspectives to generate highly representative embeddings for the target user-item pair. MacGNN (Chen et al., 2024) attempts to cluster similar micro nodes into a macro node to reduce computational complexity.
A.3. Click-Through Rate Prediction
Click-Through Rate (CTR) Prediction is one of the core algorithms in the recommendation system  (Zhu et al., 2021; Zhang et al., 2021), which predicts whether a user will click on an advertisement or not. The CTR prediction is influenced by the user’s historical click-through rate and behavior, and it is trained on vast amounts of historical data by combining various factors and characteristics. Traditional shallow CTR models, such as Logistic Regression (LR), Degree-2 Polynomial Mappings (Poly2), and Factorization Machine (FM) (Rendle, 2010), were widely used in the early days due to their lightweight nature and strong interpretability. With the development of deep learning, modern deep learning methods have achieved significant improvements. FNN uses FM in the embedding layer to generate hidden vectors as initialization values for embeddings. PNN (Qu et al., 2016) adds a product layer between the embedding layer and the MLP. By adding an attention mechanism between the embedding layer and the MLP, DIN (Zhou et al., 2018) can learn from the user’s past behavior. DIEN (Zhou et al., 2019) incorporates the ARGRU sequence model to model the evolution process of interest. DeepFM (Guo et al., 2017) replaces LR in the traditional wide part with FM, and DCN (Wang et al., 2017) further replaces FM with a cross layer network to handle feature crossover. DCNv2 (Wang et al., 2021) uses low-rank mixing to strike a balance between performance and delay, thereby achieving state-of-the-art (SOTA) results.
Appendix B EXPERIMENTAL DETAILS
B.1. Dataset Details
We adopt three publicly available datasets for offline evaluation. The statistics of these datasets are presented in Table 4. The detailed descriptions of the datasets are as follows:
  • Taobao (Zhu et al., 2018) is a user behavior dataset provided by Alibaba. Taobao is one of the most famous online shopping malls in China, the dataset includes the interaction behaviors between 28,910 users and 65,087 items from November to December 2017, including clicking, favoring, adding to cart and buying. The Taobao dataset is available on the Alibaba Tianchi page [^1] https://tianchi.aliyun.com/dataset/649.
  • QB-video (Yuan et al., 2022) is a large-scale real-world dataset obtained from the recommendation data of two content platform applications of Tencent. It has over 5 million users and 140 million interactions. We select a subset of video scenes, which mainly consists of 28,803 users and 15,540 items, including four basic behaviors: clicking, liking, sharing and following. The QB-video dataset can be obtained from the information page of the Tencent dataset [^2] https://github.com/yuangh-x/2022-NIPS-Tenrec/tree/master.
  • QK-article (Yuan et al., 2022) is a dataset in Tenrec under the article scene provided by Tencent. It records the behaviors of 52,052 users and 30,273 products, which mainly fall into four categories: reading, liking, favoring and following. The dataset can be accessed on the repository [^3] https://github.com/yuangh-x/2022-NIPS-Tenrec/tree/master.
Table 4. Statistics of the experimental datasets.
Datasets Users Items Click/Read Cart/Like Favor/Share Buy/Follow
Taobao 28,910 65,087 1,391,096 103,524 55,141 29,909
QB-video 28,803 15,540 1,503,207 14,261 1,955 1,675
QK-article 52,052 30,273 1,560,571 28,073 11,275 1,811
B.2. Baseline Details
We compare our proposed MGOE with ten relevant representative multi-task models, which can be roughly divided into two categories.
(I) Traditional Multi-task Models:
  • MMoE (Ma et al., 2018b) introduces a gating network to assign different expert networks to different tasks, allowing the model to share knowledge between different tasks while retaining the task-specific information processing ability.
  • PLE (Tang et al., 2020) extracts deep information layer by separating shared experts and task specific experts, which helps to solve the negative transfer problem and seesaw problem in multi-task learning.
  • ESMM (Ma et al., 2018a) effectively solves the problem of data sparsity and sample selection bias in CVR estimation, and improves the generalization ability of the model by training on the entire sample space.
  • AITM (Xi et al., 2021) proposes an adaptive information transfer multitask framework to models the sequence dependencies between user multi-step transitions through an adaptive information transfer module.
  • STEM (Su et al., 2024) proposes a novel share and task-specific embedding paradigm and it is equipped with an all forward task-specific backward gating network to facilitate the learning of task-specific embeddings and direct knowledge transfer across tasks.
  • MoME (Xu et al., 2024) designs an expert to share parameter efficiently and achieve resource utilization. It extracts from an identical over-parameterized base network by utilizing a binary mask learning mechanism composed of neuron-level model masking and weight-level expert masking.
(II) Graph-based Methods:
  • MoGENet (Hu et al., 2022) introduce a multi-channel graph neural network module designed to adaptively extract the most pertinent and significant information from the bipartite graph.
  • MMoCEG (Zhang et al., 2023b) addresses the expert representation degeneration problem through contrastive learning, and introduces a graph-based gating mechanism to recognize the commonalities and differences among multiple regions.
  • POGCN (Zhang et al., 2024a) introduces partial order graphs to unify separate behavior graphs into a joint graph, leveraging graph convolutional networks to aggregate neighbor information while considering behavior relations.
  • MacGNN (Chen et al., 2024) resolves the computational complexity problems in the infrastructure by reducing the node count from billions to hundreds, it improves the accuracy of CTR prediction while remaining computationally efficient.