### v4

# Macro Graph of Experts for Billion-Scale Multi-Task Recommendation

Abstract.
Graph-based multi-task learning at billion-scale presents a significant challenge, as different tasks correspond to distinct billion-scale graphs. Traditional multi-task learning methods often neglect these graph structures, relying solely on individual user and item embeddings. However, disregarding graph structures overlooks substantial potential for improving performance. In this paper, we introduce the Macro Graph of Experts (MGOE) framework, the first approach capable of leveraging macro graph embeddings to capture task-specific macro features while modeling the correlations between task-specific experts. Specifically, we propose the concept of a Macro Graph Bottom, which, for the first time, enables multi-task learning models to incorporate graph information effectively. We design the Macro Prediction Tower to dynamically integrate macro knowledge across tasks. MGOE has been deployed at scale, powering multi-task learning for a leading billion-scale recommender system, Alibaba. Extensive offline experiments conducted on three public benchmark datasets demonstrate its superiority over state-of-the-art multi-task learning methods, establishing MGOE as a breakthrough in multi-task graph-based recommendation. Furthermore, online A/B tests confirm the superiority of MGOE in billion-scale recommender systems.
multi-task learning, graph neural networks, billion-scale recommender system, click-through rate prediction
journalyear: 2026copyright: ccconference: Proceedings of the 32nd ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.1; August 09–13, 2026; Jeju Island, Republic of Koreabooktitle: Proceedings of the 32nd ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.1 (KDD ’26), August 09–13, 2026, Jeju Island, Republic of Koreadoi: 10.1145/3770854.3783958isbn: 979-8-4007-2258-5/2026/08ccs: Information systems Recommender systemsccs: Information systems Business intelligenceccs: Information systems Online advertising
1. Introduction
Multi-task learning (MTL) lies at the core of nearly all billion-scale recommendation systems (Wang et al., 2023; Caruana, 1997; Kokkinos, 2017; Zhang et al., 2023b), responsible for predicting multiple key tasks such as clicks, likes, and adding to cart (Juan et al., 2016; Zhu et al., 2023). Its goal is to address multiple tasks in real-time to meet diverse user needs. Previous multi-task recommendation models overlook the valuable information embedded within graphs. As shown in Figure 1, their embedding-based multi-task bottoms rely solely on individual user and item embeddings and disregard the rich relational semantics encoded in graph structures, which ultimately results in suboptimal performance. However, incorporating graph structures in MTL presents greater challenges compared to single-task situations due to two primary reasons: 1) online systems have billions of users and items, necessitating efficient and effective handling of this scale, and 2) multiple graph relations are formed by different tasks, requiring appropriate graph management techniques.
The mainstream approach in industrial recommendation typically employs single embeddings for multi-task problems without considering graph structures (Sener and Koltun, 2018; Ruder et al., 2019; Shen et al., 2025). ESMM (Ma et al., 2018b) enhances the robustness and improves the overall performance by utilizing the entire exposed sample space for training. MMoE (Ma et al., 2018a) pioneers the use of a basic mixture-of-experts structure, enabling task-specific predictions to learn from other tasks. PLE (Tang et al., 2020) separates task-sharing and task-specific experts to avoid the negative transfer and seesaw phenomenon. STEM (Su et al., 2024) introduces a novel shared and task-specific embeddings paradigm to effectively capture user preferences. Subsequent works focused on designing more sophisticated expert structures, such as multi-layer or task-specialized experts. Other studies explore graph-based multi-task learning; for example, MoGENet (Hu et al., 2022) proposes a multi-channel graph neural networks (GNNs) to model high-order information using user-item bipartite graphs. MMoCEG (Zhang et al., 2023a) introduces graph-based MoE to recognize the commonalities and differences among multiple regions. However, graph-based MTL models often struggle with the high computational burden of GNNs when facing billion-scale graphs.
Figure 1. Comparison of Multi-Task Learning Architectures, where ”G” denotes a gating mechanism and ”MGE” represents the proposed Macro Graph Expert. (a) Embedding-based MTL relies on feature representations. (b) Graph-based MTL utilizes GNNs to capture relational semantics. (c) Macro Graph MTL introduces macro graph bottom to incorporate macro graph information.
In general, current multi-task learning models face trade-offs in billion-scale recommender systems:
1. Lack of High-Order Information: Graph neural networks have demonstrated significant performance improvements, but single-embedding MTL approaches cannot benefit from extracting and aggregating high-order information.
2. Computational Complexity: Since each task forms a different graph, managing multiple billion-scale graphs introduces considerable computational overhead for online recommendation systems.
Considering these trade-offs, designing a billion-scale graph neural network structure that addresses multi-task learning scenarios is a promising direction. Recently, MacGNN (Chen et al., 2024) introduced a macro graph structure for single click-through rate (CTR) prediction tasks. However, applying macro structures to multi-task learning presents the following challenges:
1. Multi-task Graph Design: Designing appropriate graph structures to capture and integrate information from multiple tasks is a complex challenge.
2. Macro Expert Design: Developing experts capable of learning graph information to predict different tasks requires careful consideration and architectural innovation.
3. Inter-Macro Merging: Efficiently merging information across macro structures is crucial for effective multi-task learning.
By addressing these challenges, we introduce Macro Graph of Experts (MGOE), the first graph neural network architecture for billion-scale multi-task recommendation that avoids introducing unbearable computational complexity. We present the Macro Task Merging Graph (MTMG), which extracts multi-task information using a single macro graph. We then describe the design of macro graph experts that learn graph information to predict different tasks. Finally, we introduce the Macro Prediction Tower, which aggregates multi-level information to simultaneously predict multiple tasks.
Our contributions can be summarized as follows:
  • We pioneer the use of graph neural networks for billion-scale multi-task recommendation, solving the computational complexity challenges of GNNs in real-world applications.
  • We present the design of MTMG that incorporate multiple different graphs, fundamentally addressing the inability of existing multi-task models to utilize graph information.
  • We provide implementation details for deploying MGOE in a real-world billion-scale recommender system.
  • Extensive offline experiments conducted on three public benchmark datasets verify that MGOE outperforms both embedding-based multi-task learning models and graph-based recommendation models.  [^1] Source code is available at https://github.com/RainmannnnN/MGOE Besides, online A/B test have confirmed the superiority of MGOE in the billion-scale recommender systems.
2. Preliminaries
2.1. Notations
We use $\mathcal{U}$, $\mathcal{I}$, and $\mathcal{T}$ to denote the user set, item set, and task set, respectively. Each task $t\in\mathcal{T}$ is associated with a task-specific interaction matrix ${\mathbf{R}}^{(t)}$. For any user $u\in\mathcal{U}$ and item $i\in\mathcal{I}$, the entry $r^{(t)}_{ui}\in{\mathbf{R}}^{(t)}$ represents the interaction between $u$ and $i$ under task $t$. Specifically, $r^{(t)}_{ui}=1$ indicates that user $u$ has interacted with item $i$, while $r^{(t)}_{ui}=0$ indicates no interaction. Each task $t$ is also associated with a dataset $\mathcal{D}_{t}$, where each element is represented as $(u,i,y_{ui}^{(t)})\in\mathcal{D}_{t}$. Here, the label $y_{ui}^{(t)}\in\{0,1\}$ indicates whether user $u$ interacted (1) or did not interact (0) with item $i$ in task $t$. $r^{(t)}_{ui}$ and $y_{ui}^{(t)}$ are numerical equivalent but have representations.
2.2. Multi-Task Learning
Multi-task learning aims to develop the ability to handle multiple related but distinct recommendation tasks simultaneously, which may involve predicting whether a user will interact with an item by performing a specific behavior (e.g., clicking on a product, adding it to a cart, or making a purchase). While these tasks share foundational latent factors, including user preference structures and item attribute representations, each corresponds to distinct decision stages and motivations, exhibiting significant task-specific characteristics. The target of MTL is to learn a functions set $\mathcal{F}$, where each function $f_{t}\in\mathcal{F}$ takes user and item features as input to predict the probability $\hat{y}_{ui}^{(t)}\in[0,1]$ of potential interaction under the corresponding task, which is defined as $f_{t}(\cdot):(u,i)\rightarrow\hat{y}_{ui}^{(t)}$. Typically, MTL employs the binary cross-entropy loss as the loss function for each task $t$:
(1) $\mathcal{L}_{t}=-\sum_{(u,i,y_{ui}^{(t)})\in\mathcal{D}_{t}}\left[y_{ui}^{(t)}\cdot\log(\hat{y}_{ui}^{(t)})+(1-y_{ui}^{(t)})\cdot\log(1-\hat{y}_{ui}^{(t)})\right].$
The overall loss function $\mathcal{L}_{\text{MTL}}$ is obtained by weighted sum:
(2) $\mathcal{L}_{\text{MTL}}=\sum_{t\in\mathcal{T}}\alpha_{t}\cdot\mathcal{L}_{t},$
where parameter $\alpha_{t}$ denotes the relative importance of task $t$.
Figure 2. An illustration of the construction of Macro Task Merging Graph and the overview of Macro Graph of Experts.
2.3. MTL Bottom
Embedding-based bottom
The embedding-based MTL bottom transforms input features into fused representations. For each task, the multi-task bottom takes vector ${\bm{z}}=[{\bm{e}}_{u};\;{\bm{e}}_{i}]$ concatenated by relevant micro features from user ${\bm{e}}_{u}$ and item ${\bm{e}}_{i}$ as input, enabling further feature fusion for subsequent expert layers. Suppose the number of shared experts and task-specific experts is $N_{1}$ and $N_{2}$, the predicted probability $\hat{y}_{ui}^{(t)}$ for task $t$ can be expressed by:
(3) $h_{t}\left({\bm{z}}\right)$ $=\sum_{l=1}^{N_{1}}\operatorname{G}_{t,l}^{(1)}\left({\bm{z}}\right)\cdot\operatorname{E}_{s,l}\left({\bm{z}}\right)+\sum_{l=1}^{N_{2}}\operatorname{G}_{t,l}^{(2)}\left({\bm{z}}\right)\cdot\operatorname{E}_{t,l}\left({\bm{z}}\right);$
(4) $\hat{y}_{ui}^{(t)}$ $=\sigma\left(\operatorname{MLP}_{t}\left(h_{t}\left({\bm{z}}\right)\right)\right),$
where $\sigma$ denotes the sigmoid function, $\operatorname{E}_{s,l}$ and $\operatorname{E}_{t,l}$ represent the $l$-th shared experts and task-specific experts for task $t$, respectively. $\operatorname{G}_{t,l}^{(1)}(\cdot)$ and $\operatorname{G}_{t,l}^{(2)}(\cdot)$ denote the gating network operations that fuse the outputs from the corresponding experts. Note that the embedding-based MTL bottom focuses on micro features separately, neglecting the information induced by user-item interactions.
Graph-based bottom
The graph-based bottom captures the user-item collaborative connections through GNNs. Given the input of the $l$-th layer ${\bm{e}}^{l}_{v}$ for a certain user/item node $v$, the output of $(l+1)$-th layer ${\bm{e}}^{l+1}_{v}$ and the prediction $\hat{y}_{ui}^{(t)}$ for task $t$ are:
(5) ${\bm{e}}^{l+1}_{v}$ $=\operatorname{GNN}\left({\bm{e}}^{l}_{v},\sigma\left({\mathbf{W}}_{{\mathcal{G}}_{t}}\cdot\sum_{c\in\mathcal{N}(v)}\eta_{v\leftarrow c}\cdot{\bm{e}}^{l}_{c}\right)\right);$
(6) $\hat{y}_{ui}^{(t)}$ $=\sigma\left(\operatorname{MLP}_{t}\left([{\bm{e}}^{l+1}_{u};\;{\bm{e}}^{l+1}_{i}]\right)\right),$
where ${\mathbf{W}}_{{\mathcal{G}}_{t}}$ is the weight matrix of the task graph ${\mathcal{G}}_{t}$, $\eta_{v\leftarrow c}$ is implemented via attention, and $\mathcal{N}(v)$ denotes the set of neighbors of node $v$. When applied to billion-scale task graphs, the graph-based bottom encounters significant computational challenges.
2.4. Micro Graph
A micro graph is a fine-grained structure that models user-item interactions individually, where users and items are treated as micro nodes. Micro edges are instantiated through actual interactions (e.g., purchases) and are represented by an interaction matrix. Micro features characterize both micro nodes and edges. The trainable embeddings of micro nodes encode latent semantic information, while the entries of the interaction matrix $\mathbf{R}$} quantify the properties of micro edges. This structure captures granular user-item relationships and serves as a foundational component in GNN-based recommendation models, enabling fine-grained interaction analysis and prediction. Formally, a micro graph is denoted as $\mathcal{G}=(\mathcal{U},\mathcal{I},{\mathbf{R}})$.
3. Methodology
The overall framework of our proposed MGOE is illustrated in Figure 2. In this section, we firstly introduce the utilized MTMG. Subsequently, we present the MGOE with detailed structure.
3.1. Macro Task Merging Graph
We begin by constructing MTMG, which provides essential macro-level graph information for the subsequent MGOE modeling. The construction of MTMG primarily involves two core steps: (1) obtaining a merged interaction matrix, and (2) computing the macro nodes and constructing the macro edges in order.
Merging Interaction Matrix
In real-world scenarios, multiple tasks often coexist (e.g., users may click on and purchase an item simultaneously). These task combinations reflect the coexistence of diverse interaction patterns between users and items, indicating varying degrees of user preference. To accurately capture these preferences, we introduce a scoring function to quantify the importance level of such task combinations. For any user $u\in\mathcal{U}$ and item $i\in\mathcal{I}$, the potential mutual task combinations set is represented by:
$\mathbf{C}_{ui}=\{\left(u,i,t_{j_{1}}\right),\left(u,i,t_{j_{2}}\right),\ldots,\left(u,i,t_{j_{k}}\right)\mid t_{j_{1}},t_{j_{2}},\ldots,t_{j_{k}}\in\mathcal{T}\}.$
Let the collection of $\mathbf{C}_{ui}$ as $\mathcal{C}=\{\mathbf{C}_{ui}\mid\ u\in\mathcal{U},i\in\mathcal{I}\}$. Once we have defined the importance order of each task, we can obtain a rank function for task combination sets (Zhang et al., 2024), denoted as $\rho(\cdot):\mathcal{C}\to\mathbb{N}^{+}$. We further define a scoring function $g(\cdot):\mathcal{C}\to\mathbb{R}$, which assigns a merging preference score to each task combination, representing its relative importance. Specifically, the merging preference score $s_{ui}$ between user $u$ and item $i$ is defined as:
(7) $s_{ui}=g\left(\mathbf{C}_{ui}\right)=\begin{cases}\dfrac{[\rho(\mathbf{C}_{ui})+1]^{\beta}-1}{\beta}&,\text{if }\beta\neq 0\\ \ln[\rho(\mathbf{C}_{ui})+1]&,\text{if }\beta=0\end{cases}\,$,
where $\beta$ is a transformation parameter used to adjust the importance of $\mathbf{C}_{ui}$. Then, we arrange these values corresponding to user-item pairs to construct the merging interaction matrix $\mathbf{S}$}:
(8) ${\mathbf{S}}=[s_{ui}]_{m\times n},$
where $s_{ui}$ denotes each entry of the matrix, and $m=\left|\mathcal{U}\right|$ and $n=\left|\mathcal{I}\right|$ represent the number of users and items, respectively. The scoring function enables systematic and quantitative evaluation of task combinations, ensuring that MTMG accurately captures complex patterns in user-item interactions. This enhances both the expressiveness of the macro graph and the efficiency of MGOE in downstream MTL. Following the definition of the scoring function, we proceed to construct the nodes and edges within MTMG.
Macro Node Computation
Computing a macro node involves leveraging the merging preference score computed by Equation 7. Existing approaches treat all nodes in the macro graph uniformly, ignoring the rich semantic relationships in different task combinations. To overcome this limitation, we propose Preference-Based Grouping, a novel method that assigns distinct weights to each micro node based on its merging preference score. Specifically, we construct macro nodes from two complementary perspectives.
Let ${\mathbf{S}}^{\mathcal{U}}={\mathbf{S}}{\mathbf{S}}^{\top}$ and ${\mathbf{S}}^{\mathcal{I}}={\mathbf{S}}^{\top}{\mathbf{S}}$. Then we can obtain macro nodes of by performing clustering on users and items based on ${\mathbf{S}}^{\mathcal{U}}$ and $\mathbf{S}$$^{\mathcal{I}}$, respectively. The following takes clustering of users as an example. Let $d_{vj}$ denote the $(v,j)$-th element of matrix ${\mathbf{S}}^{\mathcal{U}}$. Subsequently, we assign different weights to different micro nodes. We define a merging preference score baseline $s_{0}$ for node $v$, such that if the sum of the merging preference score values exceeding $s_{0}$ surpasses a preset threshold $\delta_{1}$, the node is considered active, and a relatively high weight $w^{+}$ is assigned. Conversely, if the sum falls below a certain value $\delta_{2}\,\left(\delta_{2}\leq\delta_{1}\right)$, the node is regarded as inactive, and a relatively low weight $w^{-}$ is assigned. For all other cases, a weight of 1 is assigned. Mathematically, the weight assigned to micro nodes $v$ is defined as:
(9) $w_{v}=\begin{cases}w^{+}\,,&\sum_{j=1}^{m}\mathbbm{1}_{\{d_{vj}\geq s_{0}\}}\geq\delta_{1}\\ w^{-}\,,&\sum_{j=1}^{m}\mathbbm{1}_{\{d_{vj}\geq s_{0}\}}\leq\delta_{2}\\ 1,&\text{otherwise}\end{cases}\,$,
where $\mathbbm{1}_{\{\cdot\}}$ denotes the indicator function. Then, the centroid update formula of preference-based grouping can be expressed as:
(10) $\bm{\mu}_{k}=\frac{\sum_{v\in\mathcal{V}_{k}}w_{v}\cdot\mathbf{d}_{v}}{\sum_{v\in\mathcal{V}_{k}}w_{v}},$
where $\bm{\mu}_{k}$ is the centroid of cluster $k$, $\mathcal{V}_{k}$ denotes the set of micro nodes assigned to cluster $k$, $w_{v}$ is the weight associated with node $v$, and $\mathbf{d}_{v}=(d_{v1},d_{v2}\cdots,d_{vm})^{\top}$ is the $v$-th row of ${\mathbf{S}}^{\mathcal{U}}$. The clustering of items can be performed using the same method based on $\mathbf{S}$$^{\mathcal{I}}$. Preference-based grouping ensures that micro nodes with larger weights exert greater influence on the centroid’s position, thereby making the centroid determination more flexible and better aligned with users’ actual interests and task patterns.
Macro Edge Construction
Within the MTMG, macro edges illustrate the interactions between pairs of macro nodes within defined user or item subgraphs, signifying the task trends within these subgraphs. Unlike micro edges, which link predetermined user and item nodes, macro edges are designed to adaptively represent the varying strengths of connections between two macro nodes, customized for each user or item. Specifically, we formalize the entire macro nodes set in the MTMG as $\widetilde{\mathcal{V}}_{\text{MTMG}}=\left\{\mathcal{V}_{k_{1}},\mathcal{V}_{k_{2}},\ldots,\mathcal{V}_{k_{\widetilde{m}+\widetilde{n}}}\right\}$ with $\widetilde{m}$ and $\widetilde{n}$ denote the number of user clusters and item clusters, respectively. For any node $v$ and its $j^{\text{th}}$-hop neighbors, the weight of the macro edges is denoted by:
(11) $\widetilde{{\mathbf{S}}}_{v;\;\widetilde{p},\,\widetilde{q}}^{(j)}=\sum_{a\in{\mathcal{V}}_{v;\;\widetilde{p}}^{(j-1)}}\sum_{b\in{\mathcal{V}}_{v;\;\widetilde{q}}^{(j)}}s_{ab}\,$,
where $\mathcal{V}_{v;\;\widetilde{p}}^{(j-1)}$ = $\mathcal{V}_{v;\;\widetilde{p}}\ \cap\ \mathcal{N}_{v}^{(j-1)}$ denotes the macro nodes related to node $v$ within its $(j-1)^{\text{th}}$-hop neighbors and $\mathcal{V}_{v;\;\widetilde{q}}^{(j)}$ = $\mathcal{V}_{v;\;\widetilde{q}}\ \cap\ \mathcal{N}_{v}^{(j)}$ represent the macro nodes related to node $v$ within its $j^{\text{th}}$-hop neighbors. By merging different task combinations, MTMG provides a macro-level view of user-item interactions. This approach allows us to observe overarching trends and structures that might not be evident from individual interactions alone.
Macro Task Merging Graph Definition
[…]
3.2. Macro Graph Bottom
In this paper, we propose the macro graph bottom (MGB) with a mixture of macro graph experts (MGE) to exploit rich information from a macro perspective. In MGB, for each task $t$ and a certain node $v$, we assign task-specific macro embeddings to adequately capture the unique macro characteristics:
(12) $\widetilde{{\bm{e}}_{v}^{t}}=\operatorname{Lookup}\left(v,\widetilde{{\mathbf{E}}}_{t}\right),$
where $\widetilde{{\bm{e}}_{v}^{t}}$ represent the macro embeddings. In addition to its own micro features, MGB applies the tailored MGEs to extract information from macro neighbors, achieving a more comprehensive understanding of the relationships and interactions between entities in MTMG from a macro perspective. For a given user-item pair from $\mathcal{D}_{t}$ in MTMG, the input of MGB is concatenated by its micro embedding and macro neighbors embedding:
(13) ${\bm{z}}_{t}=[{\bm{e}}_{u};\;{\bm{e}}_{i};\;\widetilde{{\mathbf{E}}}_{t}(\widetilde{\mathcal{N}}_{u});\;\widetilde{{\mathbf{E}}}_{t}(\widetilde{\mathcal{N}}_{i})],$
where ${\bm{e}}_{u}$ and ${\bm{e}}_{i}$ represents the original micro features and $\widetilde{{\mathbf{E}}}_{b}(\widetilde{\mathcal{N}}_{v})$ denotes the corresponding macro embeddings of $v$’s macro neighbors, and the task $t$ output of MGB is illustrated as follow:
(14) $\operatorname{MGB}\left(\mathcal{D}_{t},\widetilde{{\mathbf{E}}},\widetilde{\mathcal{N}}\right)=\sum_{{t^{\prime}}\in\mathcal{T}}\operatorname{G}_{t^{\prime}}^{(t)}\left(\operatorname{MGE}_{t}({\bm{z}}_{t})\right)\cdot\operatorname{MGE}_{t^{\prime}}\left({\bm{z}}_{t^{\prime}}\right),$
where $\operatorname{G}_{t^{\prime}}^{(t)}(\cdot)$ denotes the gating network operation of task $t$, and $\operatorname{MGE}_{t^{\prime}}(\cdot)$ is defined as Equation 18. $\operatorname{MGB}$ integrates user and item embeddings with their corresponding macro neighbor embeddings, resulting in a more informative representation that reflects the overall macro structure.
3.3. Macro Graph Expert
Embedding-based multi-task experts, implemented using an MLP, face limitations in effectively leveraging the macro information in MTMG. We introduce a tailored MGE designed to exploit macro information efficiently. Specifically, we compute macro weights to determine the preference of the target node $v$ with respect to its macro neighbors. These macro weights are derived from the connected macro edges. Furthermore, different task patterns can cause significant disparities in macro edge weights, which we address using logarithmic smoothing. Formally, for a target node $v$ with its $j^{\text{th}}$-hop macro neighbor $\widetilde{p}$, we apply logarithmic smoothing with a parameter $\gamma$ to measure the importance of these macro neighbors. The macro weight $w_{v;\;\widetilde{p}}^{(j)}$ is computed as follows:
(15) $w_{v;\;\widetilde{p}}^{(j)}=\operatorname{softmax}\left(\log\left(\sum_{\widetilde{q}\in\widetilde{\mathcal{N}}_{v}^{(j-1)}}\widetilde{{\mathbf{S}}}_{v;\;\widetilde{p},\,\widetilde{q}}^{(j)}+1\right),\gamma\right),$
where $\gamma$ is a hyperparameter controlling the degree of smoothing (Chen et al., 2023). To aggregate macro embeddings efficiently, we employ an attention mechanism (Vaswani et al., 2017) instead of computationally expensive graph convolutional neural networks. The aggregated macro embedding $\widetilde{{\mathbf{Z}}}_{v,\widetilde{p}}$ is given by:
(16) ${\mathbf{M}}_{\text{sim}}$ $=\frac{({\mathbf{Q}}_{v}\cdot\widetilde{{\mathbf{E}}}_{\widetilde{p}})\cdot\left({\mathbf{K}}_{v}\cdot{\mathbf{E}}_{v}\right)^{\top}}{\sum_{\widetilde{n}\in\widetilde{\mathcal{N}}_{v}^{(j)}\setminus\{\widetilde{p}\}}({\mathbf{Q}}_{v}\cdot\widetilde{{\mathbf{E}}}_{\widetilde{n}})\cdot\left({\mathbf{K}}_{v}\cdot{\mathbf{E}}_{v}\right)^{\top}};$
[…]
where ${\mathbf{Q}}_{v},{\mathbf{K}}_{v}$ and ${\mathbf{V}}_{v}$ are learnable query, key, and value matrices for node $v$, and ${\mathbf{M}}_{\text{sim}}$ reflects the similarity between target node $v$ and macro nodes $\widetilde{p}$. Finally, the fusion of macro neighbor representations of task $t$ is expressed as:
(18) $\operatorname{MGE}_{t}({\bm{z}}_{t})=\sum_{\widetilde{p}\in\widetilde{\mathcal{N}}_{v}^{(j)}}w_{v,\widetilde{p}}\cdot\widetilde{{\mathbf{Z}}}_{v,\,\widetilde{p}}\,$,
where $\operatorname{MGE}_{t}({\bm{z}}_{t})$ represents the final output of macro graph expert.
3.4. Macro Prediction Tower
Macro and Micro Fusion
To capture the detailed local interactions provided by micro neighbors and the broader context offered by macro neighbors, we introduce the macro and micro fusion to aggregate information from both perspectives. Specifically, for a given target node $v$ with its numerous neighbors $\mathcal{N}_{v}$ related to task $t$, we sum its micro neighbors. Then, the micro neighbors representation ${\bm{e}}_{v;\;\text{nbr}}^{t}$ can be expressed as:
(19) ${\bm{e}}_{v;\;\text{nbr}}^{t}=\sum_{c\in\mathcal{N}_{v}}\langle{\bm{e}}^{t}_{c},{\bm{e}}_{v}^{t}\rangle\cdot{\bm{e}}^{t}_{c},$
where $\langle{\bm{e}}^{t}_{c},{\bm{e}}_{v}^{t}\rangle$ indicates the weight between the target node and its micro neighbors. Then, we mix the macro and micro informative representations through a multi-gate expert layer:
(20) ${\bm{e}}_{\text{cat}}^{t}$ $=[{\bm{e}}_{u};\;{\bm{e}}_{i};\;\operatorname{MGB}({\bm{z}}_{t});\;{\bm{e}}_{u;\;\text{nbr}}^{t};\;{\bm{e}}_{i;\;\text{nbr}}^{t}];$
(21) $x^{t}$ $=h_{t}\left({\bm{e}}_{\text{cat}}^{t}\right),$
where ${\bm{e}}_{\text{cat}}$ represents the concatenated embedding, $h_{t}(\cdot)$ is defined as Equation 4, and $x^{t}$ denotes the informative fused output from both the macro and micro levels.
To exploit possible connections between tasks and fully utilize a user’s macro preferences for sparse tasks, we replace the traditional multi-task tower with a macro correlated layer. Specifically, we use a macro cascading readout module to explore the relationships across different tasks and apply macro task adjustment to align the recommendations with the user’s macro preferences.
Macro Cascading Readout Module
Traditional multi-task recommendation models (Ma et al., 2018a; Tang et al., 2020) obtain results in a parallel manner, which ignores the possible connections between different tasks. To address this, we introduce the macro cascading readout module that generates the final output in a cascaded fashion. Specifically, the output of the previous task is fed into the next prediction tower as additional input. Suppose $\hat{y}_{ui}^{(t_{j-1})}$ is the output of the previous task, then $\hat{y}_{ui}^{(t_{j})}$ can be expressed as:
(22) $\hat{y}_{ui}^{(t_{j})}=\begin{cases}\sigma(\operatorname{MLP}(x^{t_{1}}))\,,&t_{j}=t_{1}\\ \sigma(\operatorname{MLP}(x^{t_{j}})+w_{t_{j-1}}\cdot\hat{y}_{ui}^{(t_{j-1})})\,,&t_{j}\in\mathcal{T}\setminus\{t_{1}\}\\ \end{cases}\,$,
where $w_{t_{j-1}}$ is a trainable parameter that controls the strength of learning the correlation between different tasks.
Macro Task Adjustment
With the goal of recommending more precisely to the user’s individual preferences and enhancing the model’s ability to predict sparse tasks, we propose macro task adjustment to guide the final prediction. First, based on MTMG, we construct a user macro task lookup table according to different task interactions. Specifically, we associate various user’s task interactions with the corresponding macro nodes. For a certain user $u$, the final prediction $\hat{\mathrm{Y}}_{ui}$ can be expressed as:
(23) $\hat{\mathrm{Y}}_{ui}={\mathbf{W}}_{u}\odot{\mathbf{Y}}_{\text{cat}}=\sigma(T_{u})\odot[\hat{y}_{ui}^{(t_{1})};\;\hat{y}_{ui}^{(t_{2})};\;\ldots]\,$,
where $T_{u}$ is the users’ macro task lookup table, ${\mathbf{W}}_{u}$ denotes the user-specific macro task adjustment weights. The prediction is computed via the Hadamard product (element-wise multiplication) $\odot$.
3.5. Model Training
We utilize binary cross-entropy as the loss function during the model’s training. Specifically, with the loss function $\mathcal{L}_{t}$ for a certain task $t$, the objective of MTL $\mathcal{L}_{\text{MTL}}$ is given by Equation 2. Then, the overall objective function for MGOE is formulated as:
(24) $\mathcal{L}_{\text{MGOE}}=\mathcal{L}_{\text{MTL}}+\lambda_{1}\cdot\|\mathbf{\theta}\|_{1}+\lambda_{2}\cdot\|\mathbf{\theta}\|_{2}^{2}\,$,
where $\lambda_{1}\cdot\|\mathbf{\theta}\|_{1}$ and $\lambda_{2}\cdot\|\mathbf{\theta}\|_{2}^{2}$ represent the $\ell_{1}$ and $\ell_{2}$ regularization terms, respectively, included to prevent overfitting.
4. Experiments
Table 1. Multi-task recommendation performance comparison results. The best and second-best results in each row are highlighted in bold font and underlined, respectively.
Models MMoE PLE ESMM AITM STEM MoME MoGENet MMoCEG POGCN MacGNN MGOE Improv.(%)
Taobao Click AUC 0.5580 0.5578 0.5577 0.5584 0.5561 0.5691 0.5784 0.5720 0.5841 0.5891 0.6158 4.53%
Logloss 0.3730 0.3723 0.3749 0.3922 0.3765 0.4415 0.4897 0.4134 0.3667 0.3641 0.3609 0.88%
Favor AUC 0.5672 0.5663 0.6034 0.5688 0.6192 0.6481 0.6412 0.6250 0.5693 0.5967 0.8016 23.68%
Logloss 0.1525 0.1519 0.1496 0.1628 0.1580 0.1737 0.2211 0.1607 0.1547 0.1497 0.1301 13.03%
Cart AUC 0.5459 0.5469 0.5660 0.5493 0.5666 0.5744 0.5823 0.5679 0.5714 0.5837 0.6652 13.96%
Logloss 0.2470 0.2472 0.2471 0.2666 0.2534 0.2889 0.3333 0.2679 0.2458 0.2424 0.2351 3.01%
Buy AUC 0.5560 0.5550 0.5549 0.5686 0.5676 0.5836 0.6286 0.5971 0.6142 0.6927 0.7098 2.47%
Logloss 0.0966 0.0963 0.0961 0.1012 0.1038 0.1158 0.1356 0.1012 0.0972 0.0930 0.0907 2.47%
Average AUC 0.5567 0.5565 0.5705 0.5612 0.5773 0.5937 0.6076 0.5905 0.5847 0.6155 0.6981 13.42%
Logloss 0.2172 0.2169 0.2169 0.2307 0.2229 0.2549 0.2949 0.2358 0.2161 0.2120 0.2042 3.68%
QB-video Click AUC 0.7891 0.7914 0.7934 0.7936 0.7910 0.8289 0.7786 0.8386 0.8402 0.8415 0.8474 0.70%
Logloss 0.4569 0.4560 0.4547 0.4524 0.4593 0.4337 0.4731 0.4199 0.4192 0.4191 0.4047 3.44%
Like AUC 0.6468 0.6328 0.7361 0.7547 0.7539 0.7250 0.5787 0.7665 0.6673 0.6786 0.8533 11.32%
Logloss 0.0403 0.0400 0.0350 0.0358 0.0408 0.0368 0.0727 0.0357 0.0428 0.0396 0.0300 14.29%
Share AUC 0.5125 0.5016 0.5726 0.5122 0.6403 0.5385 0.5524 0.6110 0.5270 0.6188 0.6925 8.15%
Logloss 0.0088 0.0086 0.0081 0.0167 0.0107 0.0081 0.0243 0.0133 0.0143 0.0080 0.0078 2.50%
Follow AUC 0.6128 0.6067 0.6606 0.5000 0.6792 0.6299 0.6277 0.6915 0.6336 0.6504 0.7424 7.36%
Logloss 0.0073 0.0076 0.0070 0.0131 0.0076 0.0072 0.0230 0.0080 0.0131 0.0071 0.0069 1.43%
Average AUC 0.6403 0.6331 0.6906 0.6401 0.7161 0.6805 0.6343 0.7269 0.6670 0.6973 0.7839 7.84%
Logloss 0.1283 0.1281 0.1262 0.1295 0.1296 0.1215 0.1483 0.1192 0.1224 0.1185 0.1123 5.23%
QK-article Read AUC 0.7089 0.7070 0.7124 0.7080 0.7249 0.7415 0.7732 0.7667 0.7387 0.7911 0.8006 1.20%
Logloss 0.1421 0.1514 0.1164 0.1284 0.1139 0.1077 0.1128 0.1174 0.1119 0.0894 0.0879 1.68%
Like AUC 0.7969 0.7937 0.8118 0.7964 0.8165 0.8028 0.8036 0.8338 0.7959 0.8701 0.9327 7.19%
Logloss 0.0865 0.0936 0.0721 0.0883 0.0719 0.0775 0.0815 0.0830 0.0795 0.0658 0.0519 21.12%
Favor AUC 0.8222 0.8228 0.8344 0.8226 0.8286 0.8315 0.8171 0.8516 0.7911 0.8891 0.9294 4.53%
Logloss 0.0393 0.0403 0.0323 0.0409 0.0329 0.0360 0.0389 0.0349 0.0392 0.0309 0.0256 17.15%
Follow AUC 0.7329 0.7326 0.6664 0.7427 0.7345 0.7404 0.6381 0.7620 0.7054 0.8075 0.8174 1.23%
Logloss 0.0088 0.0087 0.0077 0.0089 0.0073 0.0077 0.0157 0.0076 0.0126 0.0072 0.0070 2.78%
Average AUC 0.7652 0.7640 0.7562 0.7674 0.7761 0.7790 0.7580 0.8035 0.7578 0.8394 0.8700 3.65%
Logloss 0.0692 0.0735 0.0571 0.0666 0.0565 0.0572 0.0622 0.0607 0.0608 0.0483 0.0431 10.77%
In this section, we conduct comprehensive experiments on three real-world datasets, aiming to answer the following research questions. RQ1: How does MGOE perform compared to current state-of-the-art multi-task models? RQ2: What are the effects of the different components of MGOE on the overall performance? RQ3: In what way do the hyper-parameters in MGOE affect the overall performance? RQ4: How efficient is the proposed MGOE? RQ5: How does MGOE perform on billion-scale recommendation platforms?
4.1. Experimental Setup
4.1.1. Datasets
We conduct extensive experiments on three real-world benchmark datasets: Taobao (Zhu et al., 2018), QB-video (Yuan et al., 2022) and QK-article (Yuan et al., 2022). The detailed description and statistics of these datasets are illustrated in Appendix A.1.
4.1.2. Compared Baselines
To verify the effectiveness of MGOE, we compare with thirteen relevant representative models, which can be roughly divided into three categories. (i) Traditional Multi-task Models: MMoE, PLE, ESMM, AITM, STEM, MoME. (ii) Graph-based Methods: MoGENet, MMoCEG, POGCN, MacGNN. (iii) Single-Task Learning Methods: DeepFM, DIN, DIEN. We leave the details of these baseline models in Appendix A.2.
4.1.3. Implementation Details
All baselines and MGOE are implemented based on Pytorch. In the offline experiments, the embedding size fix to 10 and use Xavier (Glorot and Bengio, 2010) for initialization for all models. For MGOE, we set the batch size to 1024 and use Adam (Kingma, 2014) for optimization. The learning rate is searched from {1e-2, 5e-2, 1e-3, 5e-3} and the regularization term is searched from {1e-4, 5e-5, 1e-5}.
4.1.4. Evaluation Metrics
We adopt two widely used multi-task evaluation metrics, AUC (Fawcett, 2006) and Logloss (Chen et al., 2016), for a comprehensive evaluation. Higher values of AUC indicate better prediction performance of the model, while the lower values of Logloss imply better prediction accuracy. We run the experiments five times with different seeds with standard deviation to prevent extreme cases, the result is shown in Table 6 in Appendix A.3. Additionally, the results of the performance of single-task learning (STL) models are included in Table 7 in Appendix A.4.
4.2. Main Results (RQ1)
In this subsection, we compare our tailored MGOE with ten state-of-the-art baseline models on three experimental datasets. The comparison results on the AUC and Logloss metrics are reported in Table 1. From the result, we can have the following observations:
MGOE can achieve significant improvements over state-of-the-art methods across all types of tasks on all experimental datasets. From Table 1, we observe that the customized MGOE attains the highest average AUC values and the lowest average Logloss values across all task types, outperforming both traditional multi-task models and graph-based methods. Specifically, in terms of the AUC metric, MGOE surpasses the best baseline with average AUC improvements of 13.42% on Taobao, 7.84% on QB-video, and 3.65% on QK-article. Regarding the Logloss metric, MGOE achieves average effective gains of 3.68%, 5.23%, and 10.77% respectively. These comparison results validate that MGOE can produce more reasonable multi-task recommendation outcomes.
Introducing graph information can effectively improve the performance of the models. Based on the results, we can observe that graph-based models, like MMoCEG and MacGNN, generally yield better results compared to traditional multi-task models. This implies that traditional multi-task models treat users and items in isolation, overlooking the relationship between them. Conversely, integrating information from the user-item bipartite graph can mirror the more profound interactions between users and items. This is of great significance for enhancing the overall recommendation performance.
Introducing graph information from the micro level can not fully utilize the task patterns. By comparing the micro graph based methods and macro graph models in Table 1, it can be observed that macro graph models are capable of achieving relatively superior performance in multi-task recommendation compared to micro graph ones. This indicates that the traditional microscopic graph fails to effectively translate the diverse tasks of users towards items into enhanced model performance. Consequently, there is a need to integrate these user task patterns in a novel manner and consider these user and item nodes from a macroscopic perspective, which serves as the impetus for proposing MTMG.
Table 2. Ablation study results between MGOE with its three variants on Taobao and QB-video.
Variant Avg. AUC ($\uparrow$) Avg. Logloss ($\downarrow$)
Taobao MGOE 0.6936±0.0040 0.2057±0.0010
w/o preference 0.6919±0.0048 0.2060±0.0009
w/o adjustment 0.6836±0.0047 0.2080±0.0031
w/o cascading 0.6927±0.0058 0.2066±0.0009
QB-video MGOE 0.7715±0.0114 0.1124±0.0020
w/o preference 0.7700±0.0201 0.1127±0.0023
w/o adjustment 0.7513±0.0162 0.1140±0.0040
w/o cascading 0.7640±0.0135 0.1142±0.0033
4.3. Ablation Study (RQ2)
In order to verify the effectiveness of key designed components in MGOE, we conduct the ablation study and compare MGOE with its three variants: (1) w/o preference removes the preference-based grouping module when constructing MTMG, which ignores the task combination. (2) w/o adjustment excludes the macro task adjustment to overlook users’ macro interest. (3) w/o cascading, which the final output of MGOE is in a parallel fashion.
We conduct ablation studies on Taobao and QB-video datasets. The average of AUC and Logloss computed over all tasks is employed as the evaluation indicator. The results are presented in Table 2. By analyzing this table, we arrive at the following observations:
  • Distinguishing the micro nodes helps form a better macro graph. Ignoring the task relations between users and items decreases the performance of MGOE. This indicates MTMG is a more suitable macro graph for multi-task recommendation.
  • The macro prediction tower facilitates the macro information into more accurate recommendations. When the macro task adjustment is removed, the performance deteriorates, as the macro nodes preserve a wealth of users’ preference information. Removing the macro cascading readout module leads to a decline in performance. This indicates the necessity of representing the transfer of macro knowledge across different tasks.
Figure 3. Parameter study of $w_{v}$ on Taobao and QB-video.
Figure 4. Parameter study of $\gamma$ and $\widetilde{m}$ on Taobao and QB-video.
4.4. Hyper-parameter Study (RQ3)
We investigate three key hyperparameters in MGOE: (1) the weight $w_{v}$ in preference-based grouping; (2) the temperature parameter $\gamma$ for weighting macro nodes; (3) the number of macro user nodes in MTMG, denoted as $\widetilde{m}$. We conduct grid search within a small range over key hyperparameters on the Taobao and QB-video datasets. While examining $w_{v},\gamma,\widetilde{m}$, we keep all other hyperparameters fixed. The results are shown in Figure  3 and Figure 4.
Effect of Preference-Based Grouping Weight. We set the value of $w^{+}$ to either 2 or 1, and $w^{-}$ to either 1 or 0.5. Thus, we can represent the four combinations as follows: $w_{1}$=(2, 0.5), $w_{2}$=(2, 1), $w_{3}$=(1, 0.5) and $w_{4}$=(1, 1), where the first value corresponds to $w^{+}$ and the second to $w^{-}$. From the results presented in Figure 3, we can observe that in both datasets, assigning larger weights to active micro nodes and relatively smaller weights to inactive micro nodes generally leads to better performance.
Effect of Temperature Parameter. We evaluate the impact of the temperature parameter $\gamma$ over the range from 0.1 to 1.9 with an increment of 0.2. The results are shown in the top half of Figure 4. From the line chart, we can observe that, in the Taobao dataset, as $\gamma$ increases, the model performance initially improves and then declines. The performance reaches its peak when $\gamma=1.1$. In the QB-video dataset, MGOE performs best when $\gamma=1.5$, and then declines with the increase of $\gamma$.
Effect of Macro Node Number. We vary the number of macro user nodes from 5 to 30 in steps of 5. From the results presented in the bottom half of Figure 4, we can observe that an overly small number of macro user nodes tends to lead to overly coarse segmentation, resulting in poor outcomes. Additionally, selecting a relatively appropriate number of macro nodes can yield satisfactory performance. For both Taobao and QB-video, when the number of macro user nodes reaches 20, MGOE achieves the best performance.
Figure 5. Efficiency study of the model inference time.
4.5. Efficiency Study (RQ4)
Since multi-task recommendation models are typically deployed in real-time scenarios, computational efficiency is a critical metric (Wang et al., 2019a). To verify the efficiency, we compare the average inference time, model parameter size, GPU memory usage, and training time between MGOE and several representative multi-task baselines.
4.5.1. Average Inference Time
The comparison results are presented in Figure 5. From the figure, we have the following observations: (i) MGOE is nearly as efficient as the simplest expert-based model and on par with STEM in terms of efficiency. While maintaining the same inference time as these models, our model incorporates high-order graph information, which significantly enhances the performance of multi-task recommendation. (ii) MGOE has a much shorter inference time compared to graph-based models. Specifically, on the Taobao dataset, MGOE’s inference time is 55.83% faster than POGCN, 72.64% faster than MoGENet and 45.13% faster than MMoCEG. On the QK-article dataset, MGOE achieves an improvement of 56.31%, 80.49%, and 35.58% respectively. This shows that our approach to integrating high-order graph information is much more computationally efficient, enabling faster and more responsive recommendations. (iii) Among all the models, those without expert networks (MacGNN, ESMM) have the fastest inference speed, as the main source of inference time comes from the expert architecture.
Table 3. Model parameter size, GPU memory usage and training time comparison on Taobao dataset.
Models Parameter Size/M GPU Usage/MiB Training Time/s
MGOE 5.04 1194 258
MoGENet 11.57 2992 418
MMoCEG 12.22 1486 391
MMoE 3.21 1056 333
PLE 13.29 1302 957
STEM 13.59 1480 438
MacGNN 1.50 1080 320
4.5.2. Model Parameter Size
In Table 3, we provide a detailed comparison of model parameter size. Among expert-based models, MGOE demonstrates advantages with a parameter size of 5.04M. Compared to other models such as PLE (13.29M) and MMoCEG (12.22M), MGOE has a significantly reduced parameter size, specifically reduced by approximately 62.08% and 58.76% respectively. Although some models like MacGNN have a smaller parameter size (1.50M), MGOE maintain a balance between model complexity and performance, making it more suitable for real-world applications.
4.5.3. GPU Memory Usage and Training Time
Table 3 presents the GPU memory usage (in MiB) and the training time (in seconds) of MGOE and other competing models on the Taobao dataset. Specifically, MGOE requires only 1194 MiB of GPU memory, which represents a reduction of approximately 60.10% compared to MoGENet and 19.66% compared to MMoCEG. Additionally, our model’s GPU usage is also lower than that of PLE and STEM, and slightly higher than that of MMoE and MacGNN. These results demonstrate the efficiency of our model in terms of computational resource usage. Besides, the training time of MGOE is remarkably shorter compared to other methods, indicating that MGOE can converge smoothly.
4.6. Online Evaluation (RQ5)
We conducted an online A/B test on Alibaba platform. In this experiment, our model served as a rank model, replacing the existing online best-performed ranking models—MMoE and MacGNN. MacGNN mainly considers the clicking task, MMoE and MGOE consider “clicking”, “adding to cart”, “favoring”, and “buying”. Table 4 presents the average relative performance variation over two weeks for about 0.5 billion users and 1.2 billion items.
Table 4. Results of online A/B tests in the Alibaba platform.
A/B Test PCTR UCTR CVR GMV StayTime
v.s. MMoE +2.16% +1.63% + 5.88% + 16.46% +4.12%
v.s. MacGNN +0.93% +0.75% +3.69% +7.74% +3.26%
Compared to MMoE, firstly, MGOE demonstrates a performance improvement of 2.16% for PCTR, 1.63% for UCTR, 5.88% for CVR, and 16.46% for GMV, suggesting that our model boosts users’ inclination to interact with items and turn them into purchases. Secondly, the StayTime increases by 4.12%. This indicates that, despite both models considering a wide range of user actions, by incorporating their comprehensive macro interests, our model is more effective in attracting user clicks and driving business revenue.
Compared to MacGNN, which has a more narrow focus on clicking task, our model still shows favorable performance. The PCTR increased by 0.93%, the UCTR rose by 0.75%, the CVR increased by 3.69%, the GMV grew by 7.74%, and the user StayTime increased by 3.26%. This suggests that the broader consideration of user actions in MGOE can lead to better overall business outcomes, even when compared to a model specialized in a single task.
Overall, in the online A/B test on the Alibaba platform, all the results show that MTMG and the equipped MGOE are more suitable than the state-of-the-art online multi-task recommendation models.
5. Related Works
5.1. Multi-Task Learning
Multi-task learning aims to enhance the model’s generalization ability by training multiple tasks concurrently within the same model. In the early days of deep learning, the hard sharing mechanism of parameters was commonly employed. It could be applied to all hidden layers for all tasks, while maintaining task-specific output layers. The parameter soft sharing mechanism, on the other hand, allows each task to have its own model and parameters, thereby achieving better performance. To solve the issues of data sparsity (DS) and sample selection bias (SSB), ESSM (Ma et al., 2018b) simultaneously optimizes two related tasks in a sequential manner. MOE (Jacobs et al., 1991) employs a gated network to transform the original parameters shared by all samples into multiple sets of parameters, where each set is called an expert. MMOE (Ma et al., 2018a) proposes a multi-gate mixture-of-experts framework to explicitly learn to model task relationships. PLE (Tang et al., 2020) divides the model parameters into a private part and a shared part, which improves the robustness of multi-task learning and mitigates the negative transfer problem.
5.2. Graph Learning for Recommendation
In recent years, graphs have been increasingly incorporated into recommendation systems to enhance performance (Ying et al., 2018; Bei et al., 2025; Zhang et al., 2025; Huang et al., 2025). NGCF (Wang et al., 2019b) introduces higher-order connectivity into collaborative filtering, while LightGCN (He et al., 2020) further improves efficiency by removing non-linear transformations. However, the long inference time of such models limits their direct applicability to multi-task learning, motivating subsequent efforts to integrate graph structures into CTR prediction. MoGENet (Hu et al., 2022) adopts a multi-channel GNN to jointly model high-order signals on the user–item bipartite graph. MMoCEG (Zhang et al., 2023a) leverages a graph-based expert-sharing framework with contrastive learning to alleviate representation degeneration. GLSM (Sun et al., 2022) employs adaptive fusion to combine long-term and short-term behavioral information. GMT (Min et al., 2022) models heterogeneous local interactions to generate expressive embeddings for target user–item pairs. MacGNN (Chen et al., 2024) reduces computational cost by clustering similar micro-nodes into macro-nodes.
6. Conclusion
The introduction of the Macro Task Merging Graph (MTMG) and Macro Graph of Experts (MGOE) not only significantly advances the field of multi-task graph-based recommendation but also has practical implications for improving the performance of large-scale recommender systems. MTMG merges multiple billion-scale graphs into a unified macro structure, effectively consolidating complex graph data. MGOE leverages macro graph embeddings with macro graph experts, which provides a novel approach for integrating macro graph information and enables more efficient information extraction. Extensive offline experiments and online A/B tests demonstrate the superiority of our model.
Acknowledgements.
This work was supported in part by the National Natural Science Foundation of China (Grant Nos. 62272200, 62502008, and 7250129), in part by the Funding Scheme for Research and Innovation of FDCT (Grant No. 0021/2025/ITP1), and in part by the Youth Independent Innovation Science Fund of the National University of Defense Technology (Grant No. ZK25-61).
Appendix A Experimental Details
Table 5. Statistics of the experimental datasets.
Datasets Users Items Click/Read Cart/Like Favor/Share Buy/Follow
Taobao 28,910 65,087 1,391,096 103,524 55,141 29,909
QB-video 28,803 15,540 1,503,207 14,261 1,955 1,675
QK-article 52,052 30,273 1,560,571 28,073 11,275 1,811
Table 6. Multi-task recommendation performance comparison results over five trial runs.
Models MMoE PLE ESMM AITM STEM MoME MoGENet MMoCEG POGCN MacGNN MGOE Improv.(%)
Taobao Click AUC 0.5577±0.0005 0.5573±0.0006 0.5482±0.0026 0.5575±0.0002 0.5549±0.0009 0.5719±0.0040 0.5793±0.0138 0.5725±0.0049 0.5677±0.0049 0.6039±0.0009 0.6125±0.0035 1.42%
Logloss 0.3744±0.0017 0.3876±0.0016 0.3715±0.0003 0.3781±0.0032 0.3930±0.0105 0.4376±0.0176 0.4245±0.0273 0.4475±0.0791 0.3688±0.0004 0.3678±0.0011 0.3632±0.0015 1.25%
Favor AUC 0.5665±0.0013 0.6140±0.0011 0.5760±0.0099 0.5668±0.0011 0.6126±0.0026 0.6493±0.0106 0.6057±0.0365 0.6178±0.0125 0.5485±0.0108 0.6282±0.0048 0.7961±0.0033 22.61%
Logloss 0.1527±0.0006 0.1577±0.0025 0.1493±0.0004 0.1548±0.0016 0.1614±0.0036 0.1695±0.0074 0.1771±0.0094 0.1712±0.0239 0.1575±0.0038 0.1494±0.0008 0.1321±0.0015 11.52%
Cart AUC 0.5460±0.0008 0.5692±0.0010 0.5408±0.0025 0.5451±0.0008 0.5668±0.0016 0.5745±0.0047 0.5724±0.0177 0.5603±0.0044 0.5549±0.0031 0.5935±0.0026 0.6647±0.0014 12.00%
Logloss 0.2479±0.0007 0.2589±0.0018 0.2456±0.0005 0.2533±0.0028 0.2627±0.0072 0.2899±0.0054 0.2949±0.0153 0.2962±0.0501 0.2485±0.0005 0.2432±0.0007 0.2359±0.0004 3.00%
Buy AUC 0.5542±0.0014 0.5662±0.0041 0.5143±0.0027 0.5576±0.0039 0.5686±0.0010 0.5795±0.0061 0.6414±0.0135 0.5989±0.0045 0.6057±0.0033 0.6869±0.0069 0.7012±0.0077 2.08%
Logloss 0.0968±0.0001 0.0993±0.0017 0.0970±0.0002 0.0995±0.0014 0.1096±0.0039 0.1132±0.0042 0.1140±0.0047 0.1075±0.0104 0.1020±0.0018 0.0922±0.0002 0.0915±0.0005 0.76%
Average AUC 0.5561±0.0005 0.5767±0.0014 0.5448±0.0026 0.5567±0.0011 0.5757±0.0005 0.5938±0.0063 0.5997±0.0203 0.5874±0.0066 0.5692±0.0055 0.6281±0.0038 0.6936±0.0040 10.43%
Logloss 0.2179±0.0008 0.2259±0.0018 0.2159±0.0002 0.2214±0.0021 0.2317±0.0054 0.2526±0.0086 0.2526±0.0141 0.2556±0.0409 0.2192±0.0016 0.2119±0.0007 0.2057±0.0010 2.93%
QB-video Click AUC 0.7876±0.0031 0.7935±0.0009 0.7949±0.0002 0.7949±0.0003 0.7925±0.0008 0.8240±0.0044 0.6425±0.1038 0.8219±0.0220 0.8350±0.0035 0.8323±0.0040 0.8497±0.0070 1.76%
Logloss 0.4592±0.0021 0.4564±0.0008 0.4537±0.0003 0.4549±0.0011 0.4577±0.0032 0.4400±0.0045 3.7794±3.9909 0.4348±0.0176 0.4222±0.0035 0.4402±0.0105 0.4037±0.0067 4.38%
Like AUC 0.6385±0.0037 0.6279±0.0039 0.6251±0.0075 0.7518±0.0113 0.7593±0.0054 0.7396±0.0112 0.5503±0.0343 0.7396±0.0302 0.7264±0.0076 0.6953±0.0143 0.8431±0.0090 11.04%
Logloss 0.0406±0.0001 0.0380±0.0010 0.0367±0.0001 0.0361±0.0010 0.0388±0.0022 0.0342±0.0002 0.1509±0.0810 0.0359±0.0016 0.0379±0.0005 0.0404±0.0008 0.0310±0.0012 9.36%
Share AUC 0.5105±0.0041 0.5296±0.0180 0.5099±0.0017 0.5523±0.0080 0.5281±0.0562 0.5529±0.0345 0.5444±0.0095 0.5809±0.0382 0.5347±0.0075 0.5866±0.0230 0.6630±0.0189 13.02%
Logloss 0.0088±0.0001 0.0083±0.0002 0.0086±0.0001 0.0084±0.0001 0.0156±0.0023 0.0082±0.0001 0.0327±0.0055 0.0196±0.0140 0.0112±0.0005 0.0081±0.0001 0.0080±0.0001 1.23%
Follow AUC 0.6037±0.0035 0.6057±0.0069 0.5470±0.0070 0.6137±0.0035 0.6031±0.0031 0.5932±0.0368 0.5796±0.0476 0.6579±0.0200 0.6678±0.0053 0.6560±0.0211 0.7301±0.0107 9.33%
Logloss 0.0068±0.0001 0.0066±0.0001 0.0067±0.0002 0.0063±0.0001 0.0065±0.0001 0.0073±0.0002 0.0315±0.0062 0.0121±0.0093 0.0100±0.0004 0.0072±0.0001 0.0060±0.0001 4.76%
Average AUC 0.6351±0.0013 0.6392±0.0054 0.6145±0.0038 0.6781±0.0057 0.6707±0.0361 0.6774±0.0217 0.5792±0.0488 0.7001±0.0276 0.6910±0.0060 0.6926±0.0156 0.7715±0.0114 10.20%
Logloss 0.1289±0.0006 0.1273±0.0004 0.1264±0.0001 0.1264±0.0005 0.1296±0.0019 0.1224±0.0012 0.9986±1.0209 0.1256±0.0106 0.1203±0.0012 0.1240±0.0028 0.1124±0.0020 6.57%
QK-article Read AUC 0.6996±0.0008 0.7080±0.0020 0.7135±0.0034 0.7105±0.0047 0.7256±0.0006 0.7451±0.0045 0.7613±0.0073 0.7589±0.0019 0.7352±0.0028 0.7741±0.0049 0.7937±0.0078 2.53%
Logloss 0.1185±0.0013 0.1487±0.0028 0.1161±0.0005 0.1291±0.0032 0.1186±0.0029 0.1287±0.0081 0.1309±0.0182 0.1149±0.0037 0.1145±0.0024 0.1047±0.0088 0.0921±0.0060 12.03%
Like AUC 0.7768±0.0016 0.7947±0.0022 0.8107±0.0059 0.8003±0.0042 0.8184±0.0013 0.7892±0.0085 0.7322±0.0276 0.8284±0.0067 0.7763±0.0267 0.8670±0.0039 0.9250±0.0074 6.69%
Logloss 0.0782±0.0005 0.0933±0.0025 0.0722±0.0003 0.0845±0.0025 0.0741±0.0011 0.0776±0.0045 0.0905±0.0053 0.0770±0.0043 0.0811±0.0021 0.0698±0.0033 0.0532±0.0023 23.78%
Favor AUC 0.8001±0.0015 0.8231±0.0043 0.8303±0.0096 0.8227±0.0038 0.8287±0.0026 0.7996±0.0115 0.7447±0.0310 0.8480±0.0061 0.7760±0.0264 0.8835±0.0047 0.9351±0.0070 5.84%
Logloss 0.0369±0.0001 0.0402±0.0007 0.0325±0.0002 0.0395±0.0012 0.0341±0.0008 0.0343±0.0005 0.0468±0.0038 0.0354±0.0015 0.0409±0.0021 0.0342±0.0018 0.0251±0.0004 22.77%
Follow AUC 0.7040±0.0050 0.7385±0.0028 0.6936±0.0151 0.7421±0.0044 0.7304±0.0059 0.7150±0.0100 0.6295±0.0524 0.7597±0.0093 0.7072±0.0165 0.7863±0.0044 0.8225±0.0085 4.60%
Logloss 0.0078±0.0001 0.0088±0.0001 0.0076±0.0001 0.0089±0.0002 0.0074±0.0001 0.0076±0.0002 0.0193±0.0056 0.0077±0.0002 0.0154±0.0027 0.0075±0.0002 0.0066±0.0001 10.81%
Average AUC 0.7451±0.0014 0.7661±0.0015 0.7620±0.0024 0.7689±0.0038 0.7758±0.0014 0.7622±0.0086 0.7169±0.0296 0.7988±0.0060 0.7487±0.0181 0.8277±0.0045 0.8691±0.0070 5.00%
Logloss 0.0603±0.0005 0.0728±0.0011 0.0571±0.0002 0.0655±0.0002 0.0586±0.0011 0.0620±0.0033 0.0719±0.0082 0.0588±0.0024 0.0630±0.0023 0.0540±0.0035 0.0442±0.0022 18.15%
A.1. Dataset Details
We adopt three publicly available datasets for offline evaluation. The statistics of these datasets are presented in Table 5. The detailed descriptions of the datasets are as follows:
  • Taobao [^2] https://tianchi.aliyun.com/dataset/649 (Zhu et al., 2018) is a user behavior dataset provided by Alibaba, one of the most famous online shopping malls in China. It includes the activities between 28,910 users and 65,087 items from November to December 2017, including clicks, favors, carts, and buys.
  • QB-video [^3] https://github.com/yuangh-x/2022-NIPS-Tenrec/tree/master (Yuan et al., 2022) is a large-scale real-world dataset obtained from the recommendation data of two content platform applications of Tencent. We select a subset of video scenes that consists of 28,803 users and 15,540 items, including four basic behaviors: click, like, share, and follow.
  • QK-article33footnotemark: 3 (Yuan et al., 2022) is a content dataset under the article scene provided by Tencent. It records the behaviors of 52,052 users and 30,273 products, which mainly fall into four categories: reading, liking, favoring, and following.
A.2. Baseline Details
We compare our proposed MGOE with thirteen relevant representative models, which can be roughly divided into three categories.
Traditional Multi-Task Models: (i) MMoE (Ma et al., 2018a) introduces a gating network to assign different expert networks to different tasks. (ii) PLE (Tang et al., 2020) extracts deep information layer by separating shared experts and task specific experts. (iii) ESMM (Ma et al., 2018b) solves the problem of data sparsity and sample selection bias and improves generalizability by training the entire sample space. (iv) AITM (Xi et al., 2021) proposes an adaptive information transfer framework to model sequence dependencies. (v) STEM (Su et al., 2024) proposes a share and task-specific embedding paradigm to facilitate the learning of task-specific embeddings and direct knowledge transfer across tasks. (vi) MoME (Xu et al., 2024) designs an expert to share parameters efficiently and maximize resource utilization.
Graph-based Methods: (i) MoGENet (Hu et al., 2022) introduces a multi-channel GNN module to extract information from the bipartite graph. (ii) MMoCEG (Zhang et al., 2023a) introduces a graph-based gating mechanism to recognize commonalities and differences between multiple regions. (iii) POGCN (Zhang et al., 2024) introduces partial order graphs utilizing GNN to consider behavior relations. (iv) MacGNN (Chen et al., 2024) solves computational complexity problems in the infrastructure by reducing the number of nodes from billions to hundreds.
Single-Task Learning Methods: (i) DeepFM (Guo et al., 2017) combines a factorization machine and a DNN to model user-item relationships. (ii) DIN (Zhou et al., 2018) introduces an attention mechanism to capture the correlation between user’s historical behavior and candidate ads. (iii) DIEN (Zhou et al., 2019) further improves the accuracy of the CTR prediction by capturing the evolution of the user interests.
A.3. Experimental Results with Five Runs
This subsection provides detailed results of experiments where all models were run five times independently on each dataset. The reported metrics are presented as mean ± standard deviation, ensuring statistical robustness and reliability of the performance comparisons. From Table 6, we observe that MGOE also attains the highest average AUC values and the lowest average Logloss values across all task types. Specifically, in terms of the AUC metric, MGOE surpasses the best baseline with average AUC improvements of 10.43% on Taobao, 10.2% on QB-video, and 5% on QK-article. Regarding the Logloss metric, MGOE achieves average effective gains of 2.93%, 6.57%, and 18.15% respectively. These results further validate the consistency and stability of MGOE superiority across different runs.
A.4. Performance Comparison with STL Models
To further validate the effectiveness of MGOE, we conducted experiments using three STL models (DeepFM, DIN, and DIEN) across the three datasets. The results, as presented in Table 7, demonstrate that our model outperforms these STL baselines in both AUC and Logloss metrics, reinforcing its superiority.
Table 7. Comparison results with the STL models.
Models DeepFM DIN DIEN MGOE Improv.(%)
Taobao Click AUC 0.5501±0.0252 0.5632±0.0004 0.5644±0.0003 0.6125±0.0035 8.52%
Logloss 0.6280±0.0695 0.3743±0.0010 0.3753±0.0011 0.3632±0.0015 2.97%
Favor AUC 0.5810±0.0494 0.5721±0.0006 0.5861±0.0004 0.7961±0.0033 35.83%
Logloss 0.2235±0.0218 0.1527±0.0006 0.1523±0.0001 0.1321±0.0015 13.26%
Cart AUC 0.5432±0.0218 0.5660±0.0009 0.5670±0.0010 0.6647±0.0014 17.23%
Logloss 0.3705±0.0511 0.2463±0.0007 0.2462±0.0003 0.2359±0.0004 4.18%
Buy AUC 0.5655±0.0291 0.6191±0.0014 0.6187±0.0007 0.7012±0.0077 13.26%
Logloss 0.1401±0.0163 0.0942±0.0008 0.0946±0.0006 0.0915±0.0005 2.87%
Average AUC 0.5600±0.0314 0.5801±0.0005 0.5840±0.0006 0.6936±0.0040 18.77%
Logloss 0.3405±0.0397 0.2168±0.0006 0.2171±0.0005 0.2057±0.0010 5.12%
QB-video Click AUC 0.6948±0.1238 0.8010±0.0029 0.8013±0.0014 0.8497±0.0070 6.04%
Logloss 0.7279±0.4617 0.4419±0.0030 0.4427±0.0021 0.4037±0.0067 8.64%
Like AUC 0.6177±0.1341 0.6631±0.0026 0.6640±0.0021 0.8431±0.0090 26.97%
Logloss 0.1633±0.1037 0.0406±0.0001 0.0406±0.0001 0.0310±0.0012 23.65%
Share AUC 0.5651±0.0443 0.5571±0.0032 0.5612±0.0027 0.6630±0.0189 17.32%
Logloss 0.0267±0.0144 0.0084±0.0006 0.0081±0.0008 0.0080±0.0001 1.23%
Follow AUC 0.6169±0.0691 0.6130±0.0033 0.6271±0.0022 0.7301±0.0107 16.42%
Logloss 0.0234±0.0125 0.0079±0.0004 0.0074±0.0006 0.0060±0.0001 18.92%
Average AUC 0.6236±0.0928 0.6585±0.0030 0.6634±0.0021 0.7715±0.0114 16.29%
Logloss 0.2353±0.1481 0.1247±0.0011 0.1247±0.0009 0.1124±0.0020 9.86%
QK-article Read AUC 0.6013±0.1241 0.7817±0.0044 0.7810±0.0045 0.7937±0.0078 1.54%
Logloss 0.7391±0.5120 0.1012±0.0061 0.1022±0.0059 0.0921±0.0060 8.99%
Like AUC 0.6157±0.1505 0.8583±0.0080 0.8586±0.0075 0.9250±0.0074 7.73%
Logloss 0.4124±0.2750 0.0668±0.0023 0.0662±0.0021 0.0532±0.0023 19.64%
Favor AUC 0.6198±0.1447 0.8571±0.0087 0.8610±0.0082 0.9351±0.0070 8.61%
Logloss 0.1655±0.1061 0.0331±0.0021 0.0327±0.0020 0.0251±0.0004 23.24%
Follow AUC 0.6046±0.1062 0.7765±0.0105 0.7777±0.0100 0.8225±0.0085 5.76%
Logloss 0.0252±0.0138 0.0073±0.0087 0.0074±0.0085 0.0066±0.0001 9.59%
Average AUC 0.6104±0.1314 0.8184±0.0079 0.8195±0.0076 0.8691±0.0070 6.05%
Logloss 0.3356±0.2267 0.0521±0.0048 0.0521±0.0046 0.0442±0.0022 15.16%
A.5. Complexity Analysis
We further compare the complexity with the graph-based and the expert-based methods to further validate the efficiency of MGOE.
For the graph-based methods, for example, MoGENet, given the number of nodes $m+n$, the number of layers $L$ and the feature dimension $d$, the approximate complexity for feature passing in one layer is $O((m+n)d^{2})$ and the node feature transformation is $O((m+n)d)$. The overall complexity for $L$ layers is $O(L(m+n)(1+d)d)$. When $m+n$ become extremely large, this complexity significantly raised. For the expert-based methods (e.g., PLE and STEM), let $h_{e}$, $h_{s}$, $h_{g}$, and $h_{o}$ be the number of shared expert neurons, task-specific expert neurons, gate networks neurons, and tower neurons respectively. The complexity of expert layer is approximately $O(d(h_{e}+h_{g}+h_{s}))$. For a $L$ expert layers method, the complexity is $O(Ld(h_{e}+h_{g}+h_{s})+h_{e}h_{o})$ as the number of tasks/shared expert/task-specific expert is a constant. For MGOE, let $\widetilde{m}+\widetilde{n}$ be the number of macro nodes, the complexity of each macro graph expert is $O((\widetilde{m}+\widetilde{n})^{2}d)$. Since there are $T+1$ modules, the total complexity is $O((T+1)(\widetilde{m}+\widetilde{n})^{2}d)$. The macro and micro fusion are composed of CGC layers so the complexity is also $O(Ld(h_{e}+h_{g}+h_{s}))$. The overall complexities of MGOE is approximately $O((\widetilde{m}+\widetilde{n})^{2}d+Ld(h_{e}+h_{g}+h_{s})+h_{e}h_{o})$. Since the number of macro nodes $\widetilde{m}+\widetilde{n}$ is relatively small, the complexity of MGOE is comparable to that of the expert-based methods.
Overall, MGOE strikes a better balance between computational efficiency and representational power compared to them.
A.6. Pseudocode
The complete procedures of MGOE are outlined in Algorithm 1.
Input : Task set $\mathcal{T}$; Task-specific dataset $\mathcal{D}_{t}$
Output : Predicted labels
Macro Task Merging Graph
Merge interaction matrices into $\mathbf{S}$} using Equation 7;
Construct macro nodes and macro edges by Equation 10 and Equation 11;
Build macro task table $T$;
Macro Graph of Experts
Initialize model parameters $\Theta$;
for each mini-batch do
   Generate micro user/item embeddings ${\bm{e}}_{u}$, ${\bm{e}}_{i}$ and macro embeddings $\widetilde{{\mathbf{E}}}(\mathcal{\widetilde{N}}_{u})$, $\widetilde{{\mathbf{E}}}(\mathcal{\widetilde{N}}_{i})$;
 foreach $t\in\mathcal{T}$ do
      Compute macro weights $w$ and aggregated macro embeddings $\widetilde{{\mathbf{Z}}}$ via Equation 15 and Equation 16;
      Calculate macro graph expert output $\operatorname{MGE}_{t}$ by Equation 18;
      Compute micro neighbor representations $e_{\text{nbr}}^{t}$ using Equation 19;
      Concatenate micro and macro representations and pass through expert layer $h_{t}$ by Equation 20 and Equation 21;
   end foreach
  Compute final prediction $\hat{\mathrm{Y}}_{ui}$ by Equation 22 and Equation 23;
   Update model parameters $\Theta$ via $\mathcal{L}_{\text{MGOE}}$;
end for
return Predicted labels;
Algorithm 1 MGOE